### v1

# Few-shot Tuning of Foundation Models for Class-incremental Learning

Abstract
For the first time, we explore few-shot tuning of vision foundation models for class-incremental learning. Unlike existing few-shot class incremental learning (FSCIL) methods, which train an encoder on a base session to ensure forward compatibility for future continual learning, foundation models are generally trained on large unlabelled data without such considerations. This renders prior methods from traditional FSCIL incompatible for FSCIL with the foundation model. To this end, we propose Consistency-guided Asynchronous Contrastive Tuning (CoACT), a new approach to continually tune foundation models for new classes in few-shot settings. CoACT comprises three components: (i) asynchronous contrastive tuning, which learns new classes by including LoRA modules in the pre-trained encoder, while enforcing consistency between two asynchronous encoders; (ii) controlled fine-tuning, which facilitates effective tuning of a subset of the foundation model; and (iii) consistency-guided incremental tuning, which enforces additional regularization during later sessions to reduce forgetting of the learned classes. We perform an extensive study on 16 diverse datasets and demonstrate the effectiveness of CoACT, outperforming the best baseline method by 2.47% on average and with up to 12.52% on individual datasets. Additionally, CoACT shows reduced forgetting and robustness in low-shot experiments. As an added bonus, CoACT shows up to 13.5% improvement in standard FSCIL over the current SOTA on benchmark evaluations. We make our code publicly available at https://github.com/ShuvenduRoy/CoACT-FSCIL.
1 Introduction
Few-Shot Class-Incremental Learning (FSCIL) is a continual learning paradigm that involves sequentially learning a set of new classes with only a limited number of samples available per class [51; 44; 39]. This task is particularly challenging yet important as it mimics real-world scenarios where new classes may emerge over time, and collecting a large training set for these new classes can be difficult. In practice, FSCIL assumes the availability of a large base session where the number of classes and the number of samples per class are typically large [51]. The base session is subsequently followed by incremental learning sessions, which introduce new classes with a limited number of samples per class. In the traditional paradigm of FSCIL, the model is usually trained on the base session with fully supervised learning, while facilitating forward compatibility for learning future incremental classes [42; 21]. Techniques for forward compatibility during base training include allocating space in the embedding space [42] or identifying specific parameters of the encoder that it deems necessary to keep frozen while learning the new sessions [21].
Figure 1: Traditional FSCIL vs. FSCIL with foundation models.
With the rise of large foundation models with excellent generalization capabilities and impressive performances across different domains, the ability to continually learn new classes with very little data can even further enhance their adaptability and effectiveness in real-world applications. However, foundation models are generally trained on web-scale unlabeled data without any knowledge (or compatibility) for future continual learning. This in turn renders existing FSCIL solutions impractical for FSCIL with a foundation model and requires a fresh approach where we aim to tune off-the-shelf foundation models for new classes without losing previously learned knowledge (see Figure 1). As far as we are aware, enabling FSCIL by tuning large foundation models has not been explored in prior works.
Tuning foundation models also comes with its own challenges, particularly when dealing with limited labelled data. Despite advancements in parameter-efficient tuning [23; 13] and regularization techniques [38], fine-tuning of foundation models in a few-shot setting often leads to a decline in generalization due to overfitting, which can result in catastrophic forgetting in the context of continual learning.
To address this challenge, we propose Consistency-guided Asynchronous Contrastive Tuning (CoACT), a novel framework for class-incremental learning in foundation models in few-shot settings. CoACT comprises three components: (i) asynchronous contrastive tuning, (ii) controlled fine-tuning, and (iii) consistency-guided incremental tuning. Asynchronous contrastive tuning learns from the first incremental session using a novel asynchronous contrastive method that strikes a balance between adaptability to learn new classes and retaining generalizable knowledge of the pre-trained foundation model. For adapting the model to the first new session, we integrate learnable LoRA modules into the pre-trained encoder, and ensure generalization by enforcing consistency between two asynchronous encoders: a student encoder containing the learnable modules, and a teacher encoder identical to the pre-trained encoder, updated as the Exponential Moving Average (EMA) of the student. This prevents rapid change in the output distribution of the teacher, which in turn helps reduce overfitting in the student encoder while learning effective representations. In addition to training LoRA, to further enhance adaptability, we introduce controlled fine-tuning, which utilizes a two-step training protocol for the first incremental session. First, we train the newly added LoRA modules with a high Learning Rate (LR) for a certain number of epochs and then fine-tune the last few layers of the pre-trained parameters with a relatively lower LR. This also helps balance adaptability with generalizability. Finally, consistency-guided incremental tuning is a novel regularization technique that ensures effective learning of classes in the following incremental sessions while preventing forgetting of previously learned classes and preserving the generalization capabilities of the foundation model at the same time. To achieve this, we enforce consistency between the predictions of the learnable encoder in the incremental sessions and the frozen encoder from the first incremental session.
We conduct a comprehensive study on 16 diverse image recognition datasets to investigate the effectiveness of our method. The datasets include generic objects, fine-grained objects, scenes, satellite images, and texture recognition. Considering that traditional FSCIL methods are not compatible for FSCIL with foundation models, we first establish a set of baselines by building on the few-shot learning literature, namely prototype-learning [42], linear tuning, and LoRA tuning [19]. Our comprehensive experiments demonstrate that CoACT achieves a 2.47% average improvement over the best baseline method, with up to 12.79% performance gain on individual datasets (Figure 3). More importantly, CoACT exhibits reduced forgetting of already learned classes as the number of classes increases. We also evaluate the efficacy of CoACT in a traditional FSCIL setup, where the first incremental session is considerably large. Following established protocols, we report these results on CIFAR-100, CUB-200, and miniImageNet datasets. Our method achieves state-of-the-art performance (Table 3), outperforming existing approaches by 4.11%, 5.02%, and 13.47% on the three datasets, respectively. We provide detailed ablation studies showing the effectiveness of each component of our method. Overall, we make the following contributions:
  • For the first time, we explore FSCIL with foundation models to enable continuous tuning of the model to learn new classes in few-shot settings.
  • We propose CoACT, a novel method for that can effectively perform FSCIL with foundation models without losing its generalization capability or forgetting already learned classes. Our method consists of three key components that enrich our repertoire for tuning foundation models for more dynamic real-world settings.
  • Comprehensive experiments show the effectiveness of our method, achieving SOTA on both traditional FSCIL and FSCIL with foundation models. We also show reduced forgetting and effectiveness in very-low-shot settings. Extensive ablation and sensitivity studies show the effectiveness of each of our proposed components and their optimal setup. We make our code publicly available: https://github.com/ShuvenduRoy/CoACT-FSCIL.
[…]
Figure 2: Performance comparison on FSCIL with a foundation model.
[…]
Method CIFAR-100 CUB-200 miniIN
LIMIT 51.23 57.41 49.19
FACT 52.10 56.94 50.49
SAVC 53.12 62.50 57.11
SoftNet 55.33 56.75 54.68
BOT 58.75 63.75 59.57
SV-T 69.75 76.17 82.38
CPE-CLIP 80.52 64.60 82.77
CoACT (Ours) 84.62 81.19 96.24
Figure 3: Performance comparison on tradition FSCIL.
2 Related works
Few-shot class-incremental learning. Class-incremental learning is a continual learning process that focuses on continuous learning of new classes while retaining the knowledge of already learned ones [30]. In practice, machine learning models often need to learn new classes from a few labelled samples per class [51], while having no access to samples from already learned classes. This scenario has given rise to a new learning task called few-shot class-incremental learning or FSCIL [44]. The existing literature on FSCIL can be broadly categorized into two main groups: methods that continuously train both the encoder and classifier over each incremental session  [43; 6; 10; 49], and methods that keep the encoder frozen during the incremental learning sessions [52; 41; 48]. We provide a brief summary of both groups below.
As an example of the first group of FSCIL methods, MgSvF [49] employed a component-wise update strategy to ensure adapting to new classes while preserving knowledge of existing ones. The exemplar relation distillation framework [10] constructed and updated a graph of exemplar relationships to facilitate the integration of new classes. SoftNet [21] introduced a novel approach for identifying and freezing a sub-network of crucial parameters from the previous session and training the remaining parameters during incremental sessions. While such methods in the first group generally offer greater adaptability to new classes compared to the second group: (a) they require additional constraints to avoid overfitting to the new classes and thus catastrophic forgetting; (b) they are not generally feasible in the context of foundation models as encoders are quite large, resulting in strong overfitting in few-shot settings.
In the second group of FSCIL methods, FACT [50] introduced virtual prototypes to maximize class embedding separation while preserving their relative positions. SAVC [42] generated virtual classes during the base session training to maximize separability. Similarly, NC-FSCIL [47] pre-assigned optimally spaced prototypes to each base class, promoting diverse and distinct class representations. Methods in this group generally focus on ensuring forward compatibility by maintaining separability in the base classes in the learned embedding space to enable the effective integration of new classes without interfering with existing ones [52; 41; 48]. While such methods in the second group generally offer greater stability on already learned classes: (a) stability often comes at the cost of adaptability toward learning new classes; (b) these methods are generally not compatible with foundation models given lack of control over the base training of such off-the-shelf models.
Fine-tuning foundation models. A number of techniques have recently been proposed to tune foundation models without the need to re-train them from scratch. Adapter tuning [18] fine-tunes large pre-trained models to downstream tasks by inserting new learnable layers inside the pre-trained model. On the other hand, prompt-tuning [27] and prefix-tuning [28] add learnable prompts with the input embedding for learning the new task without tuning the pre-trained parameters of the model. Low-rank adapters [19; 22] have been introduced to reduce the computation cost of additional parameters. VPT [20] and AdapterFormer [4] explored parameter-efficient fine-tuning in the context of vision transformers. Nonetheless, such methods are not designed for continuous tuning of a foundation model since there are no inherent mechanisms in them to prevent loss of generalization capabilities and catastrophic forgetting.
3 Method
3.1 Problem formulation
In FSCIL, a model $\phi(x)$ is continuously trained over $T$ consecutive sessions of new classes. Each session follows an $n$-way, $k$-shot setup, where $n$ represents the number of classes in each session, and $k$ is the number of samples per class. Consequently, the training data for each session $t\in T$ can be represented as $\mathcal{D}_{train}^{t}=\{(x_{i},~{}y_{i})\}_{i=1}^{N_{t}}$, where $x_{i}$ and $y_{i}$ denote the $i^{\text{th}}$ sample and its corresponding label, and $N_{t}$ represents the number of samples in session $t$. We can express $\phi(x)=W^{T}f_{\theta}(x)$, where $f_{\theta}$ is an encoder and $W$ is a linear classifier. In the conventional setup of FSCIL, $f_{\theta}$ is generally trained on a large labelled base session ($\mathcal{D}_{train}^{0}$) while ensuring forward compatibility toward learning new classes, and remains frozen in the incremental sessions. If we assume $f_{\theta}$ to be a foundation model pre-trained on large web-scale data with no knowledge of future continual learning, standard FSCIL techniques are rendered incompatible since they require a labelled base session. In this work, we aim to design a method for tuning off-the-shelf pre-trained foundation models to continuously learn new classes in few-shot learning settings.
3.2 Consistency-guided Asynchronous Contrastive Tuning (CoACT)
In this section, we discuss the details of the three components of our proposed method: asynchronous contrastive tuning, controlled fine-tuning, and consistency-guided incremental tuning. Here, asynchronous contrastive tuning and controlled fine-tuning facilitate learning the first session (see Figure 4 (left)), while consistency-guided incremental tuning learns the remaining incremental sessions without forgetting the learned classes (see Figure 4 (right)).
Figure 4: Illustration of CoACT. (Left) Training on the first incremental session with asynchronous contrastive tuning and controlled fine-tuning. The student encoder contains learnable LoRA modules, while the teacher is identical to the foundation model but updated as the EMA of the student. Controlled fine-tuning enables the tuning of a subset of the foundation model with reduced LR after certain epochs. (Right) Consistency-guided incremental tuning enforces consistency between the learnable student and the frozen encoder from the first incremental session, providing additional regularization that prevents overfitting and forgetting.
3.2.1 Asynchronous contrastive tuning
To strike a balance between adaptability to new classes and retaining generalizable knowledge of the foundation model, we introduce asynchronous contrastive tuning as the first component in our framework. This involves fine-tuning the pre-trained model using our novel Asynchronous Contrastive Learning (ACL) approach while incorporating LoRA modules into the foumodel. Let $h_{i}=f_{\theta}^{(i)}(h_{i-1})$ be the output of $i^{\text{th}}$ layer of the pre-trained encoder, $h_{i-1}$ be the output of the $(i-1)^{\text{th}}$ hidden layer of the encoder, and $f_{\theta}^{(i)}$ be the $i^{\text{th}}$ layer of the encoder. With the new learnable LoRA layers, the output of the $i^{\text{th}}$ layer of the network can be represented as $h_{i}^{{}^{\prime}}=f_{\theta}^{(i)}(h_{i-1})+f^{(i)}_{\text{LoRA}}(h_{i-1})$, where, $f^{(i)}_{\text{LoRA}}$ is the $i^{\text{th}}$ LoRA layer added to the pre-trained encoder. For brevity, we denote the encoder with learnable LoRA layers as $f_{\theta^{{}^{\prime}}}$. We can train $f_{\theta^{{}^{\prime}}}$ on $\mathcal{D}_{train}^{1}$ to learn the first incremental session as:
$\mathcal{L}_{\text{{sup}}}=\mathcal{L}_{ce}(W^{T}f_{\theta^{{}^{\prime}}}(x),y).$ (1)
However, it has been shown in prior work that cross-entropy alone does not learn a well-separable embedding space [42] and has a higher over-fitting tendency, especially in a few-shot setting [38]. To reduce the possibility of overfitting and retain the generalization in the learnable encoder $f_{\theta^{{}^{\prime}}}$, we regularize its output distribution with a teacher encoder by maximizing their agreement in the embedding space. Here, the teacher encoder is identical to the pre-trained model (does not contain the LoRA modules) and learned through the Exponential Moving Average (EMA) of the student $f_{\theta^{{}^{\prime}}}$ as $\theta^{{}^{\prime\prime}}=m\cdot\theta^{{}^{\prime\prime}}+(1-m)\cdot\theta$, where $\theta$ and $\theta^{{}^{\prime\prime}}$ are the parameters of the student (without LoRA) and teacher encoders, and $m$ is the momentum parameter. Given that the teacher and student encoders differ in their architecture due to the addition of LoRA to the student, learning occurs asynchronously. This asynchronous encoder design and slow-moving update of the teacher through EMA ensures that the predictions from the teacher do not fluctuate. Since the teacher encoder is also initialized from the foundation model, consistency with the teacher effectively regularizes the student from overfitting.
In practice, we maximize the agreement between the embeddings of the student and teacher encoders on all the samples from each class using a supervised contrastive loss [24] as:
$\begin{split}\mathcal{L}_{\text{{ACL}}}=-\sum_{i}\frac{1}{|C_{i}|}\sum_{j\in C_{i}}\log\frac{\exp(\langle q_{i},k_{j}\rangle/\tau)}{\sum_{l\neq i}\exp(\langle q_{i},k_{l}\rangle/\tau)},\end{split}$ (2)
[…]
3.2.2 Controlled fine-tuning
To further enhance the adaptability of the model, we enable controlled fine-tuning of some of the pre-trained layers of the encoder. Since the newly added LoRA modules are randomly initialized, we begin by training only the LoRA modules of the student encoder with a higher learning rate for the initial $E_{c}$ epochs of training. This is followed by a fine-tuning stage where the last $C_{l}$ layers of the pre-trained encoder are fine-tuned with a reduced LR (scaled by a factor of $C_{f}$). We focus on fine-tuning only the last $C_{l}$ layers, as the later layers of a pre-trained model are responsible for learning domain-specific fine-grained features, whereas the earlier layers are more general and transferable to a wide range of tasks [32].
3.2.3 Consistency-guided incremental tuning
While the first two modules facilitate tuning the foundation model ($f_{\theta^{{}^{\prime}}}$) during the first session, later sessions also require the retention of previously learned classes. To facilitate this, we propose consistency-guided incremental tuning, which prevents forgetting by regularizing the output distribution of the student $f_{\theta^{{}^{\prime}}}$ when training on the incremental sessions. More specifically, we enforce consistency between the predictions of the student encoder and a frozen encoder after the first session, effectively discouraging substantial changes in the learned representations of the student. Let $f_{\theta_{\beta}}$ be the frozen encoder after the first session, and the frozen embedding of this encoder be $p_{i}=f_{\theta_{\beta}}(\mathcal{A}(x_{i}))$. We define our consistency regularizer as:
$\begin{split}\mathcal{L}_{\text{{CR}}}=-\sum_{i}\frac{1}{|C_{i}|}\sum_{j\in C_{i}}\log\frac{\exp(\langle q_{i},p_{j}\rangle/\tau)}{\sum_{l\neq i}\exp(\langle q_{i},p_{l}\rangle/\tau)}.\end{split}$ (3)
Finally, we train the model after the first session with: $\mathcal{L}_{\text{{CR}}}+\gamma\mathcal{L}_{\text{{ACL}}}+\lambda\mathcal{L}_{sup}$, where $\gamma$ and $\lambda$ controls the relative importance of the loss functions. We tune the LoRA modules and the classifier during incremental training while keeping the encoder frozen.
4 Experiments
4.1 Baselines
In the absence of available FSCIL with foundation model, we first establish a few baselines to better evaluate our proposed framework. To this end, we adopt three prominent methods on which to build these baselines. First, we use prototype learning, which has been shown to perform well for learning incremental classes in traditional FSCIL [42]. To apply this approach to a foundation model $f_{\theta}(x)$, we keep the encoder frozen during the incremental learning sessions, while the classifier ($W$) is learned with the prototypes [42] of the new classes: $W=\{w_{0}^{0},w_{1}^{0},\cdot\cdot,w_{|c^{0}|}^{0}\}\cup\{w_{0}^{T},w_{1}^{T},\cdot\cdot,w_{|c^{T}|}^{T}\}$. Here, the prototype $w_{c}^{t}$ of class $c$ from session $t$ is defined as the average of the embeddings of all samples of the class: ${w}_{c}^{t}=\frac{1}{|P_{i}|}\sum_{j\in P_{i}}{f_{\theta}({{x}}_{j})}$, where $P_{i}\stackrel{{\scriptstyle\text{def}}}{{=}}\{j:y_{j}=i\}$. Next, we consider linear tuning as a baseline and train the model in an incremental-frozen approach, where the encoder is trained only on the first session, followed by prototype learning for the remaining sessions [50]. Finally, we use parameter-efficient fine-tuning as the third baseline, where we add LoRA to different layers of the pre-trained encoder and train newly added LoRA while keeping the encoder frozen.
4.2 Experiment setup
Dataset. We comprehensively evaluate the performance of our proposed method on a diverse set of 16 datasets, including generic object detection (Caltech101 [12], CIFAR-100 [26], CUB-200 [45], MiniImageNet [40], VOC 2007 [11]), fine-grained recognition (OxfordPets [34], StanfordCars [25], Flower102 [33], Food101 [1], FGVCAircraft [29]), scene recognition (SUN397 [46], Country211 [36]), satellite-image (EuroSAT [14], Resisc-45 [5]), texture recognition (DTD [8]), and traffic sign recognition (GTSRB [17]).
Method Acc
Prototype learning 58.77$\pm$0.0
Linear tuning 58.89$\pm$0.23
LoRA 58.96$\pm$0.11
CoACT 61.43$\pm$0.12
Table 1: Overall performance over 16 datasets.
Implementation details. By default, we divide the classes into 10 (or 9; see Appendix for details) sessions with an equal number of classes and perform 10-shot continual training following FSCIL literature [50; 51]. The performance after each session is calculated as the average accuracy of all classes seen so far. Unless specified otherwise, accuracy refers to the accuracy after the last session, which is the average accuracy over all classes. We also evaluate CoACT on the traditional FSCIL setup, where the initial training is performed with a large labelled set from the same distribution as the incremental sessions, followed by tuning on the few-shot incremental sessions. We use a ViT-B/16 as the backbone for most of the experiments, while we present the detailed per-dataset results with ViT-B/32 and ViT-L/16 as well. The encoders are pre-trained on ImgeNet-21K [40]. The framework is implemented on PyTorch, and experiments are conducted on an Nvidia V100 GPU, where the training takes about 6 hours. More details on the hyper-parameters are provided in Appendix A.2.
4.3 Results
Figure 5: (left) Forgetting of learned classes. (right) Accuracy breakdown over the first session, remaining sessions, and all sessions.
Table 1 presents the average accuracy and standard deviation (over 3 runs) of our method over 16 datasets, in comparison to the aforementioned baselines. As we find from the table, CoACT achieves an average accuracy of 61.43% over the 16 datasets, outperforming others by an average of 2.47%. Among the baselines, LoRA shows the second-best accuracy with 58.96%. The other two baselines show relatively similar average accuracies to LoRA. Note that among the baselines, prototype learning with a frozen encoder is a deterministic process, thus a standard deviation of 0. Note that as we will later demonstrate in Table 3, the SOTA in the standard FSCIL setup achieve $\sim$1gains versus prior work.
Table 2: Performance of CoACT on 16 datasets and its comparison to the baseline methods.
Method
Encoder
Caltech101
CIFAR100
Country211
CUB200
DTD
EuroSat
FGVCAircraft
Food101
GTSRB
MiniImageNet
Flowers102
OxfordPets
Resisc-45
StanfordCars
SUN397
VOC 2007
Average
Pro. lear. ViT-B/16 83.47 75.58 7.42 74.10 54.72 62.20 15.51 71.09 15.53 95.31 98.09 88.10 49.83 18.90 65.68 64.72 58.77
ViT-B/32 83.90 74.29 6.79 69.32 52.83 68.91 13.50 63.57 19.84 92.76 97.09 84.57 45.30 16.85 61.29 66.64 57.34
ViT-L/16 85.10 75.04 7.23 76.31 54.33 65.48 19.03 74.66 12.24 98.05 98.82 87.64 54.85 24.69 66.58 60.16 60.01
Lin. tun. ViT-B/16 83.60 75.58 7.45 74.41 54.72 62.20 15.32 71.18 16.92 95.32 98.09 88.10 49.83 18.96 65.87 64.72 58.89
ViT-B/32 83.87 74.35 6.75 69.84 53.22 68.89 11.29 63.51 26.28 92.75 97.64 85.06 51.17 16.57 61.36 66.58 58.07
ViT-L/16 85.25 75.33 7.26 76.11 54.24 65.32 19.44 74.51 13.01 97.61 98.34 87.34 54.99 25.21 66.99 61.01 60.12
LoRA ViT-B/16 83.87 75.81 7.45 74.65 54.70 62.20 15.32 71.24 17.23 95.32 98.09 88.10 49.82 18.94 65.90 64.72 58.96
ViT-B/32 83.54 74.72 6.81 69.85 53.22 68.89 10.88 63.52 29.84 92.75 97.55 85.09 49.14 16.09 61.32 66.56 58.11
ViT-L/16 85.15 75.34 7.29 76.26 54.52 65.68 19.13 74.75 12.28 97.88 98.71 87.89 54.85 24.99 66.78 60.56 60.13
CoACT ViT-B/16 86.86 78.31 7.42 77.38 55.11 62.25 18.98 71.59 26.05 95.34 98.18 87.79 62.62 24.40 66.11 64.45 61.43
ViT-B/32 85.59 77.26 6.85 72.21 54.78 67.71 14.56 68.41 29.99 94.81 98.01 86.21 55.34 20.18 64.49 66.71 60.19
ViT-L/16 87.25 79.27 7.45 78.29 56.19 66.36 21.24 74.99 17.22 97.18 98.18 87.95 62.25 25.43 67.18 62.23 61.79
[…]
Figure 6: Performance vs. different shots.
Figure 7: Comparison to SOTA on traditional FSCIL.
Figure 8: Forgetting over incr. sessions.
Next, we report the forgetting of learned classes for each method, measured as the drop in accuracy w.r.t. the first session, in Figure 5 (left). As we observe in this figure, CoACT has the least amount of forgetting compared to the baselines with approximately 1.5% less forgetting than both LoRA and linear tuning, and 3.2% than prototype learning. We also present a breakdown of accuracies into the first incremental session, the remaining incremental sessions, and all sessions in Figure 5 (right). Here, for all methods, the first session shows particularly higher accuracy than other sessions since there is no interference (or forgetting) of other classes. While LoRA and linear tuning have higher overall accuracy than prototype tuning, the improvement mainly comes from the higher accuracy in the first session only. All baselines, however, perform relatively similarly in the remaining sessions. In contrast, CoACT shows higher improvement in both the first and remaining sessions.
To further evaluate the efficacy of CoACT, we investigate the performance with a different number of samples per class, including 1, 2, 4, 8, and 16. As seen in Figure 8, CoACT outperforms other methods in all settings and shows robustness in very-low-shot settings. The baselines vary in effectiveness with the different number of samples per class. For instance, prototype learning performs better than linear tuning and LoRA in 1-shot and 2-shot settings, while the other two perform better as the number of samples increases. These results indicate that linear tuning and LoRA do not learn very effective representations in very-low-shot settings, compared to our method and prototype learning.
Finally, in Table 2, we present per-dataset results using three different encoder backbones. With ViT-B/16 as the backbone, while CoACT shows average improvements of 2.66%, 2.54%, and 2.47% improvement over prototype learning, linear tuning, and LoRA, respectively. Notably, it shows up to a 12.79% improvement on some individual datasets such as Resisc-45. In general, we find relatively higher improvements for more challenging datasets. For instance, on the 5 datasets with the least accuracy (Country211, FGVCAircraft, GTSRB, Resisc-45, and StanfordCars), CoACT shows an average improvement of 6.14% over the best baseline method (LoRA). We observe similar improvements in the performance with the other two encoders (ViT-B/32 and ViT-L/16), where CoACT outperforms the best baseline (LoRA) by average accuracies of 2.08% and 1.67%.
To further evaluate the effectiveness of our method, we also investigate its performance in the traditional training setup of FSCIL. Following the existing literature, we present these results on the CIFAR-100, CUB-200, and miniImageNet datasets. In Table 3, we compare the performance of CoACT with prior works on the CIAFR-100 dataset, where we divide the existing methods into two groups: the first group, which trains randomly initialized models and the second group which use a pre-trained encoder. An extension of this table with more prior works is presented in the Appendix A.3. As we observe in the table, CoACT outperforms the previous SOTA without a pre-trained encoder (BOT [39]) by 25.88% and the SOTA with a pre-trained encoder (CPE-CLIP [9]) by 4.11%. For a fair comparison with existing methods, we evaluate SAVC [42] and BOT [39] with the same pre-trained encoder (ViT-B/16) as ours, where both methods show considerably lower performance compared to CoACT. Results on the CUB-200 and miniImageNet datasets are presented in Figure 8 (See Appendix A.3). For the CUB-200 dataset, the proposed method outperforms BOT and CPE-CLIP by 17.44% and 5.02%, respectively, while the improvements over these two SOTA methods on the miniImageNet dataset are 36.67% and 13.47%.
Table 3: Comparison to prior works on traditional FSCIL on CIFAR-100 with 60 base classes and 5-way 5-shot incremental setting. We reproduce existing SOTA methods (marked with *) with the same backbone as CoACT. Results on the CUB-200 and miniImageNet are presented in the Appendix.
Method Pre-trained Enc. Acc. in each session (%) ↑
0 1 2 3 4 5 6 7 8
CEC [48] 73.07 68.88 65.26 61.19 58.09 55.57 53.22 51.34 49.14
MetaFSCIL [7] 74.50 70.10 66.84 62.77 59.48 56.52 54.36 52.56 49.97
CLOM [53] 74.2 69.83 66.17 62.39 59.26 56.48 54.36 52.16 50.25
C-FSCIL [15] 77.47 72.40 67.47 63.25 59.84 56.95 54.42 52.47 50.47
LIMIT [51] 73.81 72.09 67.87 63.89 60.70 57.77 55.67 53.52 51.23
FACT [50] 74.60 72.09 67.56 63.52 61.38 58.36 56.28 54.24 52.10
SAVC [42] 79.85 73.70 69.37 65.28 61.91 59.27 57.24 54.97 53.12
SoftNet [21] 79.88 75.54 71.64 67.47 64.45 61.09 59.07 57.29 55.33
BOT [39] 80.25 77.20 75.09 70.82 67.83 64.86 62.73 60.52 58.75
SAVC* [42] ViT-B/16 82.98 75.35 74.01 71.51 70.64 69.78 68.98 67.84 66.24
BOT* [39] ViT-B/16 83.75 78.14 76.85 73.23 72.95 72.03 71.56 70.66 69.72
SV-T [35] SwinT 86.77 82.82 80.36 77.20 76.06 74.00 72.92 71.68 69.75
CPE-CLIP [9] CLIP-B/16 87.83 85.86 84.93 82.85 82.64 82.42 82.27 81.44 80.52
CoACT (Ours) ViT-B/16 90.46 88.46 88.11 86.94 86.98 86.52 86.39 86.0 84.63
Table 4: Performance on taditional FSCIL with a few-shot base session.
Method Base shots Acc. in each session (%) ↑
0 1 2 3 4 5 6 7 8
BOT [39] 500 80.25 77.20 75.09 70.82 67.83 64.86 62.73 60.52 58.75
CPE-CLIP[9] 500 87.83 85.86 84.93 82.85 82.64 82.42 82.27 81.44 80.52
CoACT (Ours) 1 59.04 57.79 58.91 58.65 59.46 59.64 60.01 60.42 59.91
4 72.9 71.11 71.39 70.75 71.24 70.95 71.2 71.16 70.57
8 76.53 74.78 74.71 73.95 74.34 73.89 73.99 73.81 73.05
32 83.87 82.0 81.61 80.24 80.27 79.64 79.56 79.25 78.35
128 83.93 83.1 82.69 82.21 81.94 81.54 81.16 80.90 80.69
Next, we investigate how CoACT compares to previous SOTA methods, namely BOT and CPE-CLIP, in terms of forgetting the already learned classes. As we find from Figure 8, CoACT shows a lower forgetting than previous SOTA methods over the incremental sessions. Specifically, CoACT shows a forgetting of 5.8% compared to the 7.3% of CPE-CLIP and 21.5% of BOT. This further shows the robustness of CoACT against forgetting during incremental learning sessions.
ACT C.F. C.I.T Acc
61.43
61.31
61.28
60.26
58.77
Table 5: Ablation study.
We also study how CoACT performs with a reduced number of samples per base class in the traditional FSCIL setup. For this experiment, we keep the incremental session setup the same (5-way 5-shot for CIFAR-100) but use only a few samples per class for the base session. The results of this study are presented in Table 4. As we find from this experiment, with only 1 sample per base class, CoACT achieves an accuracy of 59.91% in the final session, outperforming BOT with 500 samples per class. We also outperform CPE-CLIP with an accuracy of 80.69% in the final session when we use 128 samples per base class.
4.4 Ablation study
We present an ablation study on the proposed components of CoACT in Table 5, where ACT, C.F., and C.I.T stand for asynchronous contrastive tuning, controlled fine-tuning, and consistency-guided incremental tuning, respectively. Given that the asynchronous contrastive tuning component of our method could not be removed as it contains the trainable parameters, we start this study by removing controlled fine-tuning and consistency-guided incremental tuning modules individually and simultaneously. Interestingly, we observe that while individual removal of these components does not show considerable drops in performance, their concurrent application within our framework results in a significant boost in performance of 1.17% across 16 datasets. Finally, with the ablation of all three components and only training a linear classifier, we observe a 2.66% drop in performance.
Next, we present a comprehensive study on different parameters of CoACT in Table 6. In the first study (Table 6(a)), we study different alternates for LoRA, namely Adapter [18] and Prompt [27], where we observe that Adapter and Prompt do not perform as well as LoRA, with a final accuracy of 60.01% and 59.87% versus 60.15%, respectively. In the next table (Table 6(b)), we study the performance for different numbers of training epochs, where we find the best results when trained for 50 epochs. Next, in Table 6(c), we study the performance when adding different numbers of LoRA layers to the pre-trained model, where we observe that 12 blocks achieve slightly better results. In the next study (Table 6(d)), we investigate the performance of our method by adding learnable LoRA layers to both encoders, effectively incorporating synchronous training. This variant of our method shows a drop of 1.37% in the final accuracy, showing the importance of our proposed asynchronous teacher. We then study different parameters for the controlled tuning step, specifically the learning rate factor $C_{f}$, fine-tuning layers $C_{l}$, and fine-tuning start epoch $E_{c}$, in Table 6(e-g). The results from this study show the best performance when we fine-tune the pre-trained layers with a higher LR factor, effectively fine-tuning the pre-trained layers with lower LR and allowing smaller changes in the pre-trained weights. For $C_{l}$, we find the best results when we fine-tune half of the layers (6 out of 12) of the pre-trained encoder; training just the final layer or the whole network results in reduced accuracy. Moreover, the best result is obtained when starting the fine-tuning step after 10 epochs of training of the newly added LoRA layers. Lastly, In Table 6(h), we show the results for training different layers during incremental learning where we find the best results for tuning both the classifier and the LoRA layers during incremental tuning.
Module Accuracy
LoRA 60.15
Adapter 60.01
Prompt 59.87
Epochs Accuracy
10 58.98
20 59.81
50 60.15
Layers Accuracy
3 59.92
6 60.01
12 60.26
Ablation Accuracy
Asyn. enc. 60.26
Same enc. 58.89
$C_{f}$ Accuracy
1.0 59.4
0.5 60.37
0.1 61.14
$C_{l}$ Accuracy
All 60.51
Half 61.14
Last 60.76
$E_{c}$ Accuracy
0 60.64
10 61.21
25 60.62
Tuning layer Accuracy
Classifier (clf) 61.22
clf + LoRA 61.28
LoRA 60.94
Table 6: Ablation of different parameters of CoACT on FSCIL averaged across 16 datasets. Full experiments on all datasets are presented in the Appendix.
5 Conclusion
To enable few-shot class-incremental learning with pre-trained large vision models, we propose CoACT. Our method can effectively tune a foundation model to learn new classes without losing the generalization of the pre-training or forgetting previously learned classes. Extensive studies show the effectiveness of our method, achieving higher accuracy, lower forgetting, and robustness in low-shot settings. CoACT also outperforms prior SOTA in the standard FSCIL setup by a large margin. We present comprehensive experiments on different components of CoACT and make our code available to foster rapid developments in the area.
Limitations. Our study only focuses on class-incremental learning, but the notion of tuning a foundation model can be explored with other forms of continual learning, such as task-incremental learning. Additionally, a limitation of CoACT is that it comprises a few hyper-parameters, although our study shows a small sensitivity to those parameters in our study across 16 datasets.
Broader impact. Our work focuses on the few-shot tuning of foundation models with impressive generalization capability to effectively learn new classes with limited data. While our method does not have any direct negative impact, it comes with the same potential risk of being misused as any tuning method, where a model could be tuned to learn unwanted use cases. On the other hand, the proposed solution can potentially have a large positive impact, as the idea proposed in this work can be explored to update foundation models for new classes without the need to train from scratch, saving compute and other resources.
Appendix A Appendix
A.1 Datasets
Dataset Classes Sessions First Ses. class Cls/session
Caltech101 102 10 12 10
CIFAR-100 100 10 10 10
Country211 211 10 22 21
CUB-200 200 10 20 20
DTD 47 9 5 5
EuroSAT 10 10 1 1
FGVCAircraft 100 10 10 10
Food101 101 10 11 10
GTSRB 43 10 7 4
MiniImageNet 100 10 10 10
Flower102 102 10 12 10
OxfordPets 37 9 5 4
Resisc-45 45 9 5 5
StanfordCars 196 9 28 21
SUN397 397 10 45 44
VOC 2007 20 10 2 2
Table 7: Details on class splits over the continual sessions for different datasets
To comprehensively evaluate CoACT, we run the experiments on a set of 16 diverse datasets. Following the traditional setup of FSCIL [51; 50; 44], we divide each dataset into 10 sessions with equal classes where possible; otherwise, we divide them into 9 classes. In Table 7, we summarize the details of the total number of classes, number of sessions, and classes per session for all datasets. Training is conducted on a few shot samples from the training set of the corresponding dataset, while the accuracies are reported in the whole validation split of the corresponding datasets. The few-shot samples are constant across different experiments to ensure a fair comparison and are provided with our codebase.
A.2 Implementation details
We implement our framework in PyTorch and train the model using an SGD optimizer with a momentum of 0.9. The base learning rate is set to 0.1, with a batch size of 64, and the model is trained for 50 epochs for the first session and 5 epochs for the remaining sessions. A cosine LR decay scheduler is used to reduce the learning rate over the training epochs. The teacher encoder is updated with a momentum value of 0.999. For experiments with the traditional FSCIL setup, we train the model for 25 epochs with a learning rate of 0.001. All other implementation details are the same as described above. All experiments are conducted with 3 random seeds, and the reported results are averaged over the three runs.
A.3 Additional results
In Table 8, we extend the results from Table 1 to summarize the results of CoACT and its comparison to the baselines for all datasets over the continual learning sessions. In Table 9, we present a comparison to prior works on traditional FSCIL on CIFAR-100 with 60 base classes. This is the extension of Table 3 to include more prior works. In Table 10, we present the comparison to prior works on traditional FSCIL on CUB-200 with 100 base classes. Comparison to prior works on traditional FSCIL on miniImageNet with 60 base classes is presented in Table 11. Complete ablation studies across all 16 datasets are presented in Table 12. This table is an extension of Table 5 from the main paper. In Figure 9, we present a study on different parameters of asynchronous contrastive tuning, including the number of training epochs and the number of adapter layers. Next, in Figure 10, we summarize the study on different parameters of controlled fine-tuning, including learning rate, learning rate factor, encoder fine-tuning start layer, and encoder fine-tuning start epoch. Finally, in Figure 11, we show the study on incremental fine-tuning parameters, $\lambda$, and $\gamma$.
Table 8: Performance of ACT on 16 datasets over the incremental sessions, and its comparison to the baseline methods.
Dataset Method 1 2 3 4 5 6 7 8 9 10
CAL101 Prototype learning 82.29 84.55 82.63 76.35 76.43 78.77 82.79 82.82 82.99 83.47
Linear tuning 83.49 85.0 82.89 76.5 76.56 78.88 82.89 82.92 83.14 83.6
LoRA 87.57 87.33 84.76 77.58 77.13 79.42 83.26 83.23 83.4 83.87
CoACT (Ours) 93.52 93.12 90.66 82.07 82.75 84.1 86.77 86.64 86.91 86.86
CI100 Prototype learning 95.8 88.95 84.83 82.75 80.1 77.6 77.1 76.46 76.68 75.58
Linear tuning 95.87 89.02 84.88 82.78 80.11 77.6 77.11 76.47 76.69 75.58
LoRA 96.37 89.63 85.36 83.39 80.51 77.94 77.37 76.71 76.92 75.81
CoACT (Ours) 97.13 90.33 86.82 85.04 81.99 80.68 79.97 79.23 79.43 78.31
CO211 Prototype learning 22.41 16.98 13.97 11.45 10.76 9.55 8.82 8.08 7.76 7.42
Linear tuning 24.07 17.6 14.72 11.74 10.79 9.61 8.85 8.11 7.79 7.45
LoRA 24.11 17.63 14.75 11.76 10.77 9.6 8.84 8.1 7.79 7.45
CoACT (Ours) 23.64 16.43 14.05 11.43 10.56 9.43 8.75 8.02 7.72 7.42
CUB200 Prototype learning 87.17 78.1 79.92 77.19 78.77 78.79 76.74 74.93 73.55 74.1
Linear tuning 90.33 79.58 80.92 77.93 79.36 79.28 77.19 75.32 73.89 74.41
LoRA 91.83 80.62 81.54 78.25 79.6 79.55 77.46 75.58 74.16 74.65
CoACT (Ours) 94.78 82.38 83.26 80.35 81.56 81.13 79.58 78.13 77.11 77.38
DTD Prototype learning 86.0 77.0 74.0 68.62 65.0 61.67 58.57 57.0 54.72
Linear tuning 87.0 77.75 74.0 68.75 65.3 61.75 58.57 57.0 54.72
LoRA 86.17 77.17 73.94 68.62 65.07 61.64 58.55 56.98 54.7
CoACT (Ours) 88.17 80.25 75.11 69.08 65.27 62.08 58.9 57.4 55.11
EUSAT Prototype learning 100.0 97.99 87.08 84.91 75.55 72.19 65.45 64.51 60.66 62.2
Linear tuning 100.0 97.99 87.08 84.91 75.55 72.19 65.45 64.51 60.66 62.2
LoRA 100.0 97.99 87.08 84.91 75.55 72.19 65.45 64.51 60.66 62.2
CoACT (Ours) 100.0 98.1 87.4 85.05 75.7 72.51 65.56 64.5 60.74 62.25
AirCr. Prototype learning 19.77 18.34 13.5 17.48 17.71 17.13 14.91 15.88 15.28 15.51
Linear tuning 19.77 17.12 13.1 16.76 17.85 16.8 14.67 15.62 15.0 15.32
LoRA 19.52 16.84 13.0 16.69 17.79 16.75 14.67 15.62 15.0 15.32
CoACT (Ours) 28.57 15.45 12.57 19.85 20.7 21.08 19.42 19.77 19.09 18.98
Food101 Prototype learning 82.57 79.79 80.85 80.09 77.48 76.03 76.18 74.32 72.35 71.09
Linear tuning 83.55 80.25 81.18 80.4 77.73 76.21 76.33 74.43 72.44 71.18
LoRA 83.26 80.23 81.24 80.48 77.79 76.29 76.38 74.49 72.5 71.24
CoACT (Ours) 83.77 80.75 81.62 80.82 78.08 76.64 76.69 74.81 72.81 71.59
GTSRB Prototype learning 33.89 21.29 27.6 24.74 22.66 18.5 16.6 16.05 15.98 15.53
Linear tuning 40.83 26.59 31.25 26.87 24.64 20.45 18.4 17.75 17.52 16.92
LoRA 40.64 26.89 31.3 27.03 24.89 20.84 18.79 18.14 17.88 17.23
CoACT (Ours) 54.76 26.72 39.12 37.74 33.05 29.57 27.49 26.96 26.09 25.11
MiniIN Prototype learning 99.4 98.3 97.43 97.32 97.04 97.03 96.41 96.31 95.46 95.31
Linear tuning 99.4 98.3 97.43 97.32 97.04 97.03 96.41 96.32 95.47 95.32
LoRA 99.4 98.25 97.4 97.3 97.02 97.02 96.4 96.31 95.46 95.32
CoACT (Ours) 99.47 98.32 97.46 97.33 97.03 97.03 96.42 96.31 95.48 95.34
FLO102 Prototype learning 100.0 100.0 97.25 99.0 98.5 98.57 98.5 98.11 98.3 98.09
Linear tuning 100.0 100.0 97.25 99.0 98.5 98.57 98.5 98.11 98.3 98.09
LoRA 100.0 100.0 97.25 99.0 98.5 98.57 98.5 98.11 98.3 98.09
CoACT (Ours) 100.0 100.0 97.25 98.93 98.44 98.52 98.46 98.22 98.4 98.18
Pets Prototype learning 93.76 93.38 90.48 92.51 92.81 92.64 88.78 89.3 88.1
Linear tuning 93.76 93.38 90.48 92.51 92.81 92.64 88.78 89.3 88.1
LoRA 93.76 93.38 90.48 92.51 92.81 92.64 88.78 89.3 88.1
CoACT (Ours) 95.31 94.65 91.34 93.04 93.19 93.0 89.12 89.51 87.66
RES45 Prototype learning 78.84 77.07 70.43 63.98 58.49 54.5 51.73 51.71 49.83
Linear tuning 78.84 77.07 70.43 63.98 58.49 54.53 51.73 51.71 49.83
LoRA 79.21 77.56 70.66 64.02 58.54 54.53 51.72 51.73 49.82
CoACT (Ours) 97.12 89.64 80.53 75.65 70.53 66.33 64.1 64.34 62.62
Cars Prototype learning 45.67 29.35 26.89 25.17 23.69 21.65 20.19 19.94 18.9
Linear tuning 49.06 29.35 26.75 25.47 23.89 21.94 20.27 20.04 18.96
LoRA 49.09 29.63 26.9 25.47 23.91 21.95 20.28 20.09 18.94
CoACT (Ours) 61.82 38.24 35.84 33.09 31.4 28.35 25.4 25.76 24.4
SUN397 Prototype learning 86.95 83.75 78.98 75.5 72.83 70.52 68.65 66.61 65.68
Linear tuning 88.72 84.63 79.54 75.89 73.17 70.8 68.89 66.84 65.87
LoRA 89.5 84.9 79.76 76.01 73.24 70.9 68.94 66.93 65.9
CoACT (Ours) 89.82 84.75 79.91 76.08 73.36 70.99 69.01 67.08 66.11
VOC Prototype learning 98.86 94.27 92.37 90.44 88.46 81.79 81.28 65.39 64.4 64.72
Linear tuning 98.86 94.27 92.37 90.44 88.46 81.79 81.28 65.39 64.4 64.72
LoRA 98.86 94.29 92.38 90.45 88.46 81.79 81.28 65.39 64.4 64.72
CoACT (Ours) 99.62 94.7 92.45 90.48 88.35 81.75 81.31 65.06 64.1 64.45
Table 9: Comparison to prior works on traditional FSCIL on CIFAR-100 with 60 base classes and 5-way 5-shot incremental setting.
Method PT. Backbone Acc. in each session (%) ↑
0 1 2 3 4 5 6 7 8
Rebalancing [16] 61.31 47.80 39.31 31.91 25.68 21.35 18.67 17.24 14.17
iCaRL [37] 61.31 46.32 42.94 37.63 30.49 24.00 20.89 18.80 17.21
TOPIC [43] 61.31 50.09 45.17 41.16 37.48 35.52 32.19 29.46 24.42
IDLVQ-C [3] 64.77 59.87 55.93 52.62 49.88 47.55 44.83 43.14 41.84
FSLL [31] 66.48 61.75 58.16 54.16 51.10 48.53 46.54 44.20 42.28
FSLL+SS [31] 68.85 63.14 59.24 55.23 52.24 49.65 47.74 45.23 43.92
F2M [41] 67.28 63.80 60.38 57.06 54.08 51.39 48.82 46.58 44.65
CEC [48] 73.07 68.88 65.26 61.19 58.09 55.57 53.22 51.34 49.14
MetaFSCIL [7] 74.50 70.10 66.84 62.77 59.48 56.52 54.36 52.56 49.97
CLOM [53] 74.2 69.83 66.17 62.39 59.26 56.48 54.36 52.16 50.25
C-FSCIL [15] 77.47 72.40 67.47 63.25 59.84 56.95 54.42 52.47 50.47
LIMIT [51] 73.81 72.09 67.87 63.89 60.70 57.77 55.67 53.52 51.23
FACT [50] 74.60 72.09 67.56 63.52 61.38 58.36 56.28 54.24 52.10
SAVC [42] 79.85 73.70 69.37 65.28 61.91 59.27 57.24 54.97 53.12
SoftNet [21] 79.88 75.54 71.64 67.47 64.45 61.09 59.07 57.29 55.33
BOT [39] 80.25 77.20 75.09 70.82 67.83 64.86 62.73 60.52 58.75
SV-T [35] SwinT 86.77 82.82 80.36 77.20 76.06 74.00 72.92 71.68 69.75
CPE-CLIP[9] CLIP-B/16 87.83 85.86 84.93 82.85 82.64 82.42 82.27 81.44 80.52
Ours ViT-B/16 90.46 88.46 88.11 86.94 86.98 86.52 86.39 86.0 84.63
Table 10: Comparison to prior works across the base and incremental sessions on CUB-200.
Method Acc. in each session (%) ↑
0 1 2 3 4 5 6 7 8 9 10
iCaRL [37] 68.68 52.65 48.61 44.16 36.62 29.52 27.83 26.26 24.01 23.89 21.16
EEIL [2] 68.68 53.63 47.91 44.20 36.30 27.46 25.93 24.70 23.95 24.13 22.11
TOPIC [43] 68.68 62.49 54.81 49.99 45.25 41.40 38.35 35.36 32.22 28.31 26.28
Rebalancing [16] 68.68 57.12 44.21 28.78 26.71 25.66 24.62 21.52 20.12 20.06 19.87
SPPR [52] 68.68 61.85 57.43 52.68 50.19 46.88 44.65 43.07 40.17 39.63 37.33
MetaFSCIL[7] 75.90 72.41 68.78 64.78 62.96 59.99 58.30 56.85 54.78 53.82 52.64
F2M [41] 81.07 78.16 75.57 72.89 70.86 68.17 67.01 65.26 63.36 61.76 60.26
CEC [48] 75.85 71.94 68.50 63.50 62.43 58.27 57.73 55.81 54.83 53.52 52.28
FACT [50] 75.90 73.23 70.84 66.13 65.56 62.15 61.74 59.83 58.41 57.89 56.94
LIMIT [51] 75.89 73.55 71.99 68.14 67.42 63.61 62.40 61.35 59.91 58.66 57.41
SoftNet [21] 78.07 74.58 71.37 67.54 65.37 62.60 61.07 59.37 57.53 57.21 56.75
SAVC [42] 81.85 77.92 74.95 70.21 69.96 67.02 66.16 65.30 63.84 63.15 62.50
BOT [39] 82.31 78.03 75.45 70.99 71.06 67.85 67.44 66.05 64.95 64.31 63.75
CPE-CLIP [9] 81.58 78.52 76.68 71.86 71.52 70.23 67.66 66.52 65.09 64.47 64.60
SV-T [35] 84.19 82.63 81.21 78.97 79.38 77.64 77.55 75.71 75.91 75.77 76.17
Ours 88.68 86.26 85.83 83.38 83.52 81.71 81.77 81.77 81.02 80.76 81.19
Table 11: Comparison to prior works across the base and incremental sessions on miniImageNet.
Method Acc. in each session (%) ↑
0 1 2 3 4 5 6 7 8
iCaRL [37] 71.77 61.85 58.12 54.60 51.49 48.47 45.90 44.19 42.71
Rebalancing [16] 72.30 66.37 61.00 56.93 53.31 49.93 46.47 44.13 42.19
TOPIC [43] 61.31 50.09 45.17 41.16 37.48 35.52 32.19 29.46 24.42
EEIL [2] 61.31 46.58 44.00 37.29 33.14 27.12 24.10 21.57 19.58
FSLL [31] 66.48 61.75 58.16 54.16 51.10 48.53 46.54 44.20 42.28
FSLL+SS [31] 68.85 63.14 59.24 55.23 52.24 49.65 47.74 45.23 43.92
F2M [41] 72.05 67.47 63.16 59.70 56.71 53.77 51.11 49.21 47.84
CEC [48] 72.00 66.83 62.97 59.43 56.70 53.73 51.19 49.24 47.63
MetaFSCIL [7] 72.04 67.94 63.77 60.29 57.58 55.16 52.90 50.79 49.19
C-FSCIL [15] 76.40 71.14 66.46 63.29 60.42 57.46 54.78 53.11 51.41
FACT [50] 72.56 69.63 66.38 62.77 60.60 57.33 54.34 52.16 50.49
CLOM [53] 73.08 68.09 64.16 60.41 57.41 54.29 51.54 49.37 48.00
LIMIT [51] 72.32 68.47 64.30 60.78 57.95 55.07 52.70 50.72 49.19
SoftNet [21] 79.77 75.08 70.59 66.93 64.00 61.00 57.81 55.81 54.68
SAVC [42] 81.12 76.14 72.43 68.92 66.48 62.95 59.92 58.39 57.11
BOT [39] 84.30 79.59 75.49 71.4 68.45 65.12 62.20 60.52 59.57
CPE-CLIP [9] 90.23 89.56 87.42 86.80 86.51 85.08 83.43 83.38 82.77
SV-T [35] 88.75 87.92 86.07 84.84 84.30 83.24 82.22 82.28 82.38
Ours 97.63 97.55 97.09 97.02 97.0 96.58 96.3 96.29 96.24
Table 12: Ablation study on all datasets.
Asy. Cont.
Cal. Tun.
Doub. Con.
Caltech101
CIFAR100
Country211
CUB200
DTD
EuroSat
FGVCAircraft
Food101
GTSRB
MiniImageNet
Flowers102
OxfordPets
Resisc-45
StanfordCars
SUN397
VOC 2007
Average
86.86 78.31 7.42 77.38 55.11 62.25 18.98 71.59 26.05 95.34 98.18 87.79 62.62 24.4 66.11 64.45 61.43
85.89 78.45 7.45 77.27 54.59 62.91 18.20 71.83 26.29 95.04 98.01 88.25 61.77 24.14 66.01 64.89 61.31
86.01 78.96 7.46 77.2 54.82 62.31 18.91 71.47 26.05 95.34 98.15 87.92 61.31 24.07 65.98 64.59 61.28
84.62 78.37 7.49 76.44 54.83 62.22 18.23 71.39 18.02 95.34 98.09 88.1 57.21 23.18 65.94 64.72 60.26
83.47 75.58 7.42 74.1 54.72 62.2 15.51 71.09 15.53 95.31 98.09 88.1 49.83 18.9 65.68 64.72 58.77
Figure 9: Study on different parameters of asynchronous contrastive tuning, including the number of training epochs and number of adapter layers
Figure 10: Study on different parameters of controlled fine-tuning, including learning rate, learning rate factor, encoder fine-tuning start layer, and encoder fine-tuning start epoch.
Figure 11: Study on incremental fine-tuning parameters, $\lambda$, and $\gamma$.

### v2

@@
-Few-shot Tuning of Foundation Models for Class-incremental Learning
+Consistency-Guided Asynchronous Contrastive Tuning for Few-Shot Class-Incremental Tuning of Foundation Models
 Abstract
-For the first time, we explore few-shot tuning of vision foundation models for class-incremental learning.
-Unlike existing few-shot class incremental learning (FSCIL) methods, which train an encoder on a base session to ensure forward compatibility for future continual learning, foundation models are generally trained on large unlabelled data without such considerations.
-This renders prior methods from traditional FSCIL incompatible for FSCIL with the foundation model.
-To this end, we propose Consistency-guided Asynchronous Contrastive Tuning (CoACT), a new approach to continually tune foundation models for new classes in few-shot settings.
-CoACT comprises three components: (i) asynchronous contrastive tuning, which learns new classes by including LoRA modules in the pre-trained encoder, while enforcing consistency between two asynchronous encoders; (ii) controlled fine-tuning, which facilitates effective tuning of a subset of the foundation model; and (iii) consistency-guided incremental tuning, which enforces additional regularization during later sessions to reduce forgetting of the learned classes.
-We perform an extensive study on 16 diverse datasets and demonstrate the effectiveness of CoACT, outperforming the best baseline method by 2.47% on average and with up to 12.52% on individual datasets.
-Additionally, CoACT shows reduced forgetting and robustness in low-shot experiments.
-As an added bonus, CoACT shows up to 13.5% improvement in standard FSCIL over the current SOTA on benchmark evaluations.
+We propose Consistency-guided Asynchronous Contrastive Tuning (CoACT), a novel method for continuously tuning foundation models to learn new classes in few-shot settings.
+CoACT consists of three key components:(i) asynchronous contrastive tuning, which learns new classes by including LoRA modules in the pre-trained encoder while enforcing consistency between two asynchronous encoders; (ii) controlled fine-tuning, which facilitates effective tuning of a subset of the foundation model; and (iii) consistency-guided incremental tuning, which enforces additional regularization during later sessions to reduce forgetting of the learned classes.
+We evaluate our proposed solution on Few-Shot Class-Incremental Learning (FSCIL) as well as a new and more challenging setup called Few-Shot Class-Incremental Tuning (FSCIT), which facilitates the continual tuning of vision foundation models to learn new classes with only a few samples per class.
+Unlike traditional FSCIL, FSCIT does not require a large in-distribution base session for initial fully supervised training prior to the incremental few-shot sessions.
+We conduct extensive evaluations across 16 diverse datasets, demonstrating the effectiveness of CoACT in both FSCIL and FSCIT setups.
+CoACT outperforms existing methods by up to 5.02% in FSCIL and up to 12.51% in FSCIT for individual datasets, with an average improvement of 2.47%.
+Furthermore, CoACT exhibits reduced forgetting and enhanced robustness in low-shot experiments.
+Detailed ablation and sensitivity studies highlight the contribution of each component of CoACT.
 We make our code publicly available at https://github.com/ShuvenduRoy/CoACT-FSCIL.
 1 Introduction
-Few-Shot Class-Incremental Learning (FSCIL) is a continual learning paradigm that involves sequentially learning a set of new classes with only a limited number of samples available per class [51; 44; 39].
-This task is particularly challenging yet important as it mimics real-world scenarios where new classes may emerge over time, and collecting a large training set for these new classes can be difficult.
-In practice, FSCIL assumes the availability of a large base session where the number of classes and the number of samples per class are typically large [51].
-The base session is subsequently followed by incremental learning sessions, which introduce new classes with a limited number of samples per class.
-In the traditional paradigm of FSCIL, the model is usually trained on the base session with fully supervised learning, while facilitating forward compatibility for learning future incremental classes [42; 21].
-Techniques for forward compatibility during base training include allocating space in the embedding space [42] or identifying specific parameters of the encoder that it deems necessary to keep frozen while learning the new sessions [21].
-Figure 1: Traditional FSCIL vs. FSCIL with foundation models.
-With the rise of large foundation models with excellent generalization capabilities and impressive performances across different domains, the ability to continually learn new classes with very little data can even further enhance their adaptability and effectiveness in real-world applications.
-However, foundation models are generally trained on web-scale unlabeled data without any knowledge (or compatibility) for future continual learning.
-This in turn renders existing FSCIL solutions impractical for FSCIL with a foundation model and requires a fresh approach where we aim to tune off-the-shelf foundation models for new classes without losing previously learned knowledge (see Figure 1).
-As far as we are aware, enabling FSCIL by tuning large foundation models has not been explored in prior works.
-Tuning foundation models also comes with its own challenges, particularly when dealing with limited labelled data.
-Despite advancements in parameter-efficient tuning [23; 13] and regularization techniques [38], fine-tuning of foundation models in a few-shot setting often leads to a decline in generalization due to overfitting, which can result in catastrophic forgetting in the context of continual learning.
-To address this challenge, we propose Consistency-guided Asynchronous Contrastive Tuning (CoACT), a novel framework for class-incremental learning in foundation models in few-shot settings.
-CoACT comprises three components: (i) asynchronous contrastive tuning, (ii) controlled fine-tuning, and (iii) consistency-guided incremental tuning.
-Asynchronous contrastive tuning learns from the first incremental session using a novel asynchronous contrastive method that strikes a balance between adaptability to learn new classes and retaining generalizable knowledge of the pre-trained foundation model.
-For adapting the model to the first new session, we integrate learnable LoRA modules into the pre-trained encoder, and ensure generalization by enforcing consistency between two asynchronous encoders: a student encoder containing the learnable modules, and a teacher encoder identical to the pre-trained encoder, updated as the Exponential Moving Average (EMA) of the student.
-This prevents rapid change in the output distribution of the teacher, which in turn helps reduce overfitting in the student encoder while learning effective representations.
-In addition to training LoRA, to further enhance adaptability, we introduce controlled fine-tuning, which utilizes a two-step training protocol for the first incremental session.
-First, we train the newly added LoRA modules with a high Learning Rate (LR) for a certain number of epochs and then fine-tune the last few layers of the pre-trained parameters with a relatively lower LR.
-This also helps balance adaptability with generalizability.
-Finally, consistency-guided incremental tuning is a novel regularization technique that ensures effective learning of classes in the following incremental sessions while preventing forgetting of previously learned classes and preserving the generalization capabilities of the foundation model at the same time.
-To achieve this, we enforce consistency between the predictions of the learnable encoder in the incremental sessions and the frozen encoder from the first incremental session.
-We conduct a comprehensive study on 16 diverse image recognition datasets to investigate the effectiveness of our method.
-The datasets include generic objects, fine-grained objects, scenes, satellite images, and texture recognition.
-Considering that traditional FSCIL methods are not compatible for FSCIL with foundation models, we first establish a set of baselines by building on the few-shot learning literature, namely prototype-learning [42], linear tuning, and LoRA tuning [19].
-Our comprehensive experiments demonstrate that CoACT achieves a 2.47% average improvement over the best baseline method, with up to 12.79% performance gain on individual datasets (Figure 3).
-More importantly, CoACT exhibits reduced forgetting of already learned classes as the number of classes increases.
-We also evaluate the efficacy of CoACT in a traditional FSCIL setup, where the first incremental session is considerably large.
-Following established protocols, we report these results on CIFAR-100, CUB-200, and miniImageNet datasets.
-Our method achieves state-of-the-art performance (Table 3), outperforming existing approaches by 4.11%, 5.02%, and 13.47% on the three datasets, respectively.
-We provide detailed ablation studies showing the effectiveness of each component of our method.
-Overall, we make the following contributions:
-  • For the first time, we explore FSCIL with foundation models to enable continuous tuning of the model to learn new classes in few-shot settings.
-  • We propose CoACT, a novel method for that can effectively perform FSCIL with foundation models without losing its generalization capability or forgetting already learned classes.
-Our method consists of three key components that enrich our repertoire for tuning foundation models for more dynamic real-world settings.
-  • Comprehensive experiments show the effectiveness of our method, achieving SOTA on both traditional FSCIL and FSCIL with foundation models.
-We also show reduced forgetting and effectiveness in very-low-shot settings.
-Extensive ablation and sensitivity studies show the effectiveness of each of our proposed components and their optimal setup.
-We make our code publicly available: https://github.com/ShuvenduRoy/CoACT-FSCIL.
+Large foundation models pre-trained on web-scale unlabeled data exhibit strong generalization capabilities on downstream tasks when fine-tuned with relatively small amounts of labelled data (Radford et al., 2021; Zhou et al., 2022c).
+However, the immense size of these pre-trained models introduces significant challenges for fine-tuning, especially when working with limited labelled data (Khattak et al., 2023; Yao et al., 2023).
+Despite recent advancements such as parameter-efficient tuning (Khattak et al., 2023; Gao et al., 2023), it has been observed that fine-tuning the model in a few-shot setting often leads to a decline in the out-of-the-box generalization capability of the foundation model (Roy & Etemad, 2024).
+Yet, real-world applications not only necessitate learning from a few samples but also demand continual learning of new tasks.
+Few-shot class-incremental learning (FSCIL) (Masana et al., 2022) is a continual learning setup that extends the scope of few-shot learning by enabling models to adapt incrementally to new tasks while preserving existing knowledge.
+While few prior works (D’Alessandro et al., 2023; Park et al., 2024) have explored tuning a foundation model in FSCIL setups, existing approaches often struggle with the forgetting of leaned classes and a loss of the foundation model’s generalization capabilities.
 […]
-Figure 2: Performance comparison on FSCIL with a foundation model.
+Figure 1: Performance comparison on FSCIT with a foundation model.
 […]
@@
 CPE-CLIP 80.52 64.60 82.77
-CoACT (Ours) 84.62 81.19 96.24
-Figure 3: Performance comparison on tradition FSCIL.
+PriViLege 86.06 75.08 94.10
+CoACT (Ours) 84.63 81.19 96.24
+Figure 2: Performance comparison to existing methods on tradition FSCIL.
+In this work, we propose Consistency-guided Asynchronous Contrastive Tuning (CoACT), a novel framework for class-incremental tuning of vision foundation models in few-shot settings.
+CoACT introduces three novel components: (i) asynchronous contrastive tuning, (ii) controlled fine-tuning, and (iii) consistency-guided incremental tuning.
+To strike a balance between adaptability to learn new classes and retaining generalizable knowledge of the pre-trained foundation model asynchronous contrastive tuning learns from the first incremental session using a novel asynchronous contrastive loss.
+Specifically, to provide adaptability to learn new classes, we integrate learnable LoRA modules into the pre-trained encoder and ensure generalization by enforcing consistency between two asynchronous encoders: a student encoder containing the learnable modules and a teacher encoder identical to the pre-trained encoder, updated as the Exponential Moving Average (EMA) of the student.
+This prevents rapid change in the output distribution of the teacher, which in turn helps reduce overfitting in the student encoder while learning new classes.
+To further enhance adaptability, we introduce controlled fine-tuning, which is a two-step training protocol for training the first incremental session.
+First, we train the newly added LoRA modules with a high Learning Rate (LR) for a certain number of epochs and then fine-tune the last few layers of the pre-trained parameters with a relatively lower LR.
+This also helps balance adaptability with generalizability.
+Finally, to ensure effective learning of classes in the following incremental sessions while preventing forgetting of previously learned classes and preserving the generalization capabilities of the foundation model at the same time, we introduce a novel regularization technique, consistency-guided incremental tuning.
+We achieve this by enforcing consistency between the predictions of the learnable encoder in the incremental sessions and the frozen encoder from the first incremental session.
+We conduct a comprehensive study on 16 diverse image recognition datasets to investigate the effectiveness of our method in FSCIL.
+Additionally, we introduce a novel and more challenging setup called few-shot class-incremental tuning (FSCIT), where, unlike FSCIL, we do not assume the availability of a fully-supervised base session.
+Specifically, in FSCIT, a foundation model is fine-tuned over incremental sessions with a few samples per class for all sessions, including the first session.
+The datasets include generic objects, fine-grained objects, scenes, satellite images, and texture recognition.
+Our comprehensive experiments demonstrate that on the FSCIL setup, CoACT outperforms prior methods by up to 5.02%.
+On our proposed FSCIT setup, CoACT outperforms existing methods by 2.47% on average across 16 datasets, with up to 12.79% performance gain on individual datasets.
+More importantly, CoACT exhibits reduced forgetting of already learned classes as the number of classes increases.
+CoACT also exhibit generalization with different sizes of pre-trained encoders and different numbers of shots.
+We provide detailed ablation studies showing the effectiveness of each component of our method.
+Overall, our contributions are:
+  • We propose CoACT, which consists of three novel components to tune a vision foundation model in a few-shot continual learning setup without losing its generalization capability or forgetting the already learned classes.
+  • We introduce a new paradigm of continual learning called Few-shot Class-Incremental Tuning (FSCIT), which aims to tune a foundation model to continuously learn new classes with few samples per class.
+We also establish several baselines for this new paradigm.
+  • Comprehensive experiments show the effectiveness of our method, achieving state-of-the-art (SOTA) on both FSCIT and FSCIL setup.
+We also show reduced forgetting and effectiveness in very-low-shot settings.
+Extensive ablation and sensitivity studies show the effectiveness of each of our proposed components.
 2 Related works
-Few-shot class-incremental learning.
-Class-incremental learning is a continual learning process that focuses on continuous learning of new classes while retaining the knowledge of already learned ones [30].
-In practice, machine learning models often need to learn new classes from a few labelled samples per class [51], while having no access to samples from already learned classes.
-This scenario has given rise to a new learning task called few-shot class-incremental learning or FSCIL [44].
-The existing literature on FSCIL can be broadly categorized into two main groups: methods that continuously train both the encoder and classifier over each incremental session  [43; 6; 10; 49], and methods that keep the encoder frozen during the incremental learning sessions [52; 41; 48].
-We provide a brief summary of both groups below.
-As an example of the first group of FSCIL methods, MgSvF [49] employed a component-wise update strategy to ensure adapting to new classes while preserving knowledge of existing ones.
-The exemplar relation distillation framework [10] constructed and updated a graph of exemplar relationships to facilitate the integration of new classes.
-SoftNet [21] introduced a novel approach for identifying and freezing a sub-network of crucial parameters from the previous session and training the remaining parameters during incremental sessions.
-While such methods in the first group generally offer greater adaptability to new classes compared to the second group: (a) they require additional constraints to avoid overfitting to the new classes and thus catastrophic forgetting; (b) they are not generally feasible in the context of foundation models as encoders are quite large, resulting in strong overfitting in few-shot settings.
-In the second group of FSCIL methods, FACT [50] introduced virtual prototypes to maximize class embedding separation while preserving their relative positions.
-SAVC [42] generated virtual classes during the base session training to maximize separability.
-Similarly, NC-FSCIL [47] pre-assigned optimally spaced prototypes to each base class, promoting diverse and distinct class representations.
-Methods in this group generally focus on ensuring forward compatibility by maintaining separability in the base classes in the learned embedding space to enable the effective integration of new classes without interfering with existing ones [52; 41; 48].
-While such methods in the second group generally offer greater stability on already learned classes: (a) stability often comes at the cost of adaptability toward learning new classes; (b) these methods are generally not compatible with foundation models given lack of control over the base training of such off-the-shelf models.
-Fine-tuning foundation models.
-A number of techniques have recently been proposed to tune foundation models without the need to re-train them from scratch.
-Adapter tuning [18] fine-tunes large pre-trained models to downstream tasks by inserting new learnable layers inside the pre-trained model.
-On the other hand, prompt-tuning [27] and prefix-tuning [28] add learnable prompts with the input embedding for learning the new task without tuning the pre-trained parameters of the model.
-Low-rank adapters [19; 22] have been introduced to reduce the computation cost of additional parameters.
-VPT [20] and AdapterFormer [4] explored parameter-efficient fine-tuning in the context of vision transformers.
-Nonetheless, such methods are not designed for continuous tuning of a foundation model since there are no inherent mechanisms in them to prevent loss of generalization capabilities and catastrophic forgetting.
+2.1 Fine-tuning foundation models.
+A number of techniques have recently been proposed to tune foundation models without the need to re-train them from scratch, addressing the growing demand for computationally efficient and parameter-efficient fine-tuning methods.
+These approaches leverage the pre-trained capabilities of large models, adapting them to downstream tasks without modifying the core pre-trained parameters extensively.
+Adapter tuning (Houlsby et al., 2019) fine-tunes large pre-trained models to downstream tasks by inserting new learnable layers inside the pre-trained model.
+On the other hand, prompt-tuning (Lester et al., 2021) and prefix-tuning (Li & Liang, 2021) add learnable prompts with the input embedding for learning the new task without tuning the pre-trained parameters of the model.
+Low-rank adapters (Hu et al., 2021; Karimi Mahabadi et al., 2021) have been introduced to reduce the computation cost of additional parameters.
+VPT (Jia et al., 2022) and AdapterFormer (Chen et al., 2022) explored parameter-efficient fine-tuning in the context of vision transformers.
+A well-known issue of training a foundation model with few labelled samples per class is the overfitting and the loss of the generalization of the foundation model.
+More recent, CoPrompt Roy & Etemad (2024) proposed a novel regularization technique by enforcing consistency between the pre-trained and learnable encoders to reduce overfitting and improve generalization.
+Nonetheless, existing methods for tuning foundation models are not designed for continuous tuning of the model since there are no inherent mechanisms to prevent loss of generalization and catastrophic forgetting.
+2.2 Few-shot class-incremental learning.
+Class-incremental learning is a continual learning process that focuses on continuous learning of new classes while retaining the knowledge of already learned ones (Wang et al., 2023; Smith et al., 2023; Wang et al., 2024; 2022).
+In practice, machine learning models often need to learn new classes from a few labelled samples per class (Zhou et al., 2022b), while having no access to samples from already learned classes.
+This scenario has given rise to a new learning task called few-shot class-incremental learning or FSCIL (Tian et al., 2023; Tao et al., 2020).
+The existing literature on FSCIL can be broadly categorized into two main groups: methods that continuously train both the encoder and classifier over each incremental session  (Cheraghian et al., 2021; Dong et al., 2021; Zhao et al., 2021), and methods that keep the encoder frozen during the incremental learning sessions (Zhu et al., 2021; Shi et al., 2021; Zhang et al., 2021).
+As an example of the first group of methods, MgSvF (Zhao et al., 2021) employed a component-wise update strategy to ensure adapting to new classes while preserving knowledge of existing ones.
+The exemplar relation distillation framework (Dong et al., 2021) constructed and updated a graph of exemplar relationships to facilitate the integration of new classes.
+SoftNet (Kang et al., 2023) introduced a novel approach for identifying and freezing a sub-network of crucial parameters from the previous session and training the remaining parameters during incremental sessions.
+While such methods in the first group generally offer greater adaptability, the methods in the second group focus on ensuring stability by maintaining separability in the base classes in the learned embedding space.
+For example, FACT (Zhou et al., 2022a) introduced virtual prototypes to maximize class embedding separation while preserving their relative positions.
+SAVC (Song et al., 2023) generated virtual classes during the base session training to maximize separability.
+Similarly, NC-FSCIL (Yang et al., 2022) pre-assigned optimally spaced prototypes to each base class, promoting diverse and distinct class representations.
+Some of the more recent works have explored tuning a foundation in the FSCIL setup.
+For instance, CPE-CLIP (D’Alessandro et al., 2023) utilized the strong generalization capability of a pre-trained vision-language model to learn new classes with parameter-efficient fine-tuning.
+SV-T (Qiu et al., 2023) used a semantic-visual guided Transformer to enable few-shot class-incremental learning while reducing overfitting by combining text-based semantic labels with visual labels to improve generalization to new classes.
+PriViLege (Park et al., 2024) showed that with prompt tuning and knowledge distillation, pre-trained vision language shows strong performances in FSCIL.
+Nonetheless, these methods operate on the FSCIL setting and still show forgetting issues as they learn new classes.
 3 Method
 3.1 Problem formulation
-In FSCIL, a model $\phi(x)$ is continuously trained over $T$ consecutive sessions of new classes.
+In FSCIT, a model $\phi(x)$ is continuously trained over $T$ consecutive sessions of new classes.
 Each session follows an $n$-way, $k$-shot setup, where $n$ represents the number of classes in each session, and $k$ is the number of samples per class.
@@
 We can express $\phi(x)=W^{T}f_{\theta}(x)$, where $f_{\theta}$ is an encoder and $W$ is a linear classifier.
-In the conventional setup of FSCIL, $f_{\theta}$ is generally trained on a large labelled base session ($\mathcal{D}_{train}^{0}$) while ensuring forward compatibility toward learning new classes, and remains frozen in the incremental sessions.
-If we assume $f_{\theta}$ to be a foundation model pre-trained on large web-scale data with no knowledge of future continual learning, standard FSCIL techniques are rendered incompatible since they require a labelled base session.
-In this work, we aim to design a method for tuning off-the-shelf pre-trained foundation models to continuously learn new classes in few-shot learning settings.
-3.2 Consistency-guided Asynchronous Contrastive Tuning (CoACT)
-In this section, we discuss the details of the three components of our proposed method: asynchronous contrastive tuning, controlled fine-tuning, and consistency-guided incremental tuning.
-Here, asynchronous contrastive tuning and controlled fine-tuning facilitate learning the first session (see Figure 4 (left)), while consistency-guided incremental tuning learns the remaining incremental sessions without forgetting the learned classes (see Figure 4 (right)).
+In the conventional setup of FSCIL, $f_{\theta}$ is trained on a large labelled in-distribution base session ($\mathcal{D}_{train}^{0}$) in a fully-supervised manner.
+In FSCIT, we do not assume the availability of such a base session, rendering standard FSCIL techniques incompatible.
+In this work, we aim to design a method for tuning off-the-shelf vision foundation models to continuously learn new classes in few-shot settings.
 Figure 3: Illustration of CoACT.
@@
 Controlled fine-tuning enables the tuning of a subset of the foundation model with reduced LR after certain epochs.
-(Right) Consistency-guided incremental tuning enforces consistency between the learnable student and the frozen encoder from the first incremental session, providing additional regularization that prevents overfitting and forgetting.
-3.2.1 Asynchronous contrastive tuning
-To strike a balance between adaptability to new classes and retaining generalizable knowledge of the foundation model, we introduce asynchronous contrastive tuning as the first component in our framework.
-This involves fine-tuning the pre-trained model using our novel Asynchronous Contrastive Learning (ACL) approach while incorporating LoRA modules into the foumodel.
+(Right) Consistency-guided incremental tuning enforces consistency between the learnable student and the frozen encoder from the first session, providing additional regularization that prevents overfitting and forgetting.
+3.2 Consistency-guided Asynchronous Contrastive Tuning (CoACT)
+In this section, we discuss the details of the three components of our proposed method: asynchronous contrastive tuning, controlled fine-tuning, and consistency-guided incremental tuning.
+Here, asynchronous contrastive tuning and controlled fine-tuning facilitate learning the first session (see Figure 3 (left)), while consistency-guided incremental tuning learns the remaining incremental sessions without forgetting the learned classes (see Figure 3 (right)).
+3.2.1 Asynchronous contrastive tuning.
+To strike a balance between adaptability to learn new classes and retaining generalizable knowledge of the foundation model, we introduce asynchronous contrastive tuning as the first component in our framework.
+This involves fine-tuning the pre-trained model using our novel Asynchronous Contrastive Learning (ACL) approach while incorporating LoRA modules into the model.
 Let $h_{i}=f_{\theta}^{(i)}(h_{i-1})$ be the output of $i^{\text{th}}$ layer of the pre-trained encoder, $h_{i-1}$ be the output of the $(i-1)^{\text{th}}$ hidden layer of the encoder, and $f_{\theta}^{(i)}$ be the $i^{\text{th}}$ layer of the encoder.
@@
 For brevity, we denote the encoder with learnable LoRA layers as $f_{\theta^{{}^{\prime}}}$.
-We can train $f_{\theta^{{}^{\prime}}}$ on $\mathcal{D}_{train}^{1}$ to learn the first incremental session as:
+We can train $f_{\theta^{{}^{\prime}}}$ on $\mathcal{D}_{train}^{1}$ to learn the first session:
 $\mathcal{L}_{\text{{sup}}}=\mathcal{L}_{ce}(W^{T}f_{\theta^{{}^{\prime}}}(x),y).$ (1)
-However, it has been shown in prior work that cross-entropy alone does not learn a well-separable embedding space [42] and has a higher over-fitting tendency, especially in a few-shot setting [38].
-To reduce the possibility of overfitting and retain the generalization in the learnable encoder $f_{\theta^{{}^{\prime}}}$, we regularize its output distribution with a teacher encoder by maximizing their agreement in the embedding space.
-Here, the teacher encoder is identical to the pre-trained model (does not contain the LoRA modules) and learned through the Exponential Moving Average (EMA) of the student $f_{\theta^{{}^{\prime}}}$ as $\theta^{{}^{\prime\prime}}=m\cdot\theta^{{}^{\prime\prime}}+(1-m)\cdot\theta$, where $\theta$ and $\theta^{{}^{\prime\prime}}$ are the parameters of the student (without LoRA) and teacher encoders, and $m$ is the momentum parameter.
+However, it has been shown in prior work that cross-entropy alone does not learn a well-separable embedding space (Song et al., 2023) and has a higher over-fitting tendency, especially in a few-shot setting (Roy & Etemad, 2024).
+To reduce the possibility of overfitting and retain the generalization in the learnable encoder $f_{\theta^{{}^{\prime}}}$, we regularize its output distribution with an asynchronous teacher encoder by maximizing their agreement in the embedding space.
+Our novelty lies in the asynchronous design of the encoders, where the student encoder contains the learnable LoRA modules, but the teacher encoder is identical to the pre-trained model (without LoRA modules) and learned through the Exponential Moving Average (EMA) of the student $f_{\theta^{{}^{\prime}}}$ as $\theta^{{}^{\prime\prime}}=m\cdot\theta^{{}^{\prime\prime}}+(1-m)\cdot\theta$, where $\theta$ and $\theta^{{}^{\prime\prime}}$ are the parameters of the student (excluding LoRA) and teacher encoders, and $m$ is the momentum parameter.
+Since the teacher lacks LoRA modules, the EMA update applies only to the student’s base parameters, affecting the teacher layers corresponding to those fine-tuned in the student (Section 3.2.2).
 Given that the teacher and student encoders differ in their architecture due to the addition of LoRA to the student, learning occurs asynchronously.
@@
 Since the teacher encoder is also initialized from the foundation model, consistency with the teacher effectively regularizes the student from overfitting.
-In practice, we maximize the agreement between the embeddings of the student and teacher encoders on all the samples from each class using a supervised contrastive loss [24] as:
+In practice, we maximize the agreement between the embeddings of the student and teacher encoders on all the samples from each class as:
 $\begin{split}\mathcal{L}_{\text{{ACL}}}=-\sum_{i}\frac{1}{|C_{i}|}\sum_{j\in C_{i}}\log\frac{\exp(\langle q_{i},k_{j}\rangle/\tau)}{\sum_{l\neq i}\exp(\langle q_{i},k_{l}\rangle/\tau)},\end{split}$ (2)
 […]
-3.2.2 Controlled fine-tuning
-To further enhance the adaptability of the model, we enable controlled fine-tuning of some of the pre-trained layers of the encoder.
-Since the newly added LoRA modules are randomly initialized, we begin by training only the LoRA modules of the student encoder with a higher learning rate for the initial $E_{c}$ epochs of training.
+The exclusion of LoRA adapters from the EMA updates prevents rapid fluctuations in the teacher encoder, ensuring it serves as a slow-moving reference that mitigates overfitting and retains the generalization capabilities of the pre-trained foundation model.
+If LoRA adapters were included, the teacher would change more dynamically, reducing its ability to regularize the student effectively and undermining the balance between adaptability and generalization retention.
+Maintaining the teacher closer to the pre-trained model prevents the student from drifting too far and acts as a form of regularization that constrains learned representations while preserving prior knowledge.
+3.2.2 Controlled fine-tuning.
+Controlled fine-tuning is a two-step training protocol to enhance the adaptability of the model by selectively fine-tuning a few layers of the pre-trained encoder.
+Since the newly added LoRA modules are randomly initialized, we begin by training only the LoRA modules of the student encoder with a higher learning rate for the initial $E_{c}$ epochs of training to prevent the propagation of randomness to the well-trained pre-trained encoder, which could otherwise lead to forgetting of its generalizable knowledge.
 This is followed by a fine-tuning stage where the last $C_{l}$ layers of the pre-trained encoder are fine-tuned with a reduced LR (scaled by a factor of $C_{f}$).
-We focus on fine-tuning only the last $C_{l}$ layers, as the later layers of a pre-trained model are responsible for learning domain-specific fine-grained features, whereas the earlier layers are more general and transferable to a wide range of tasks [32].
-3.2.3 Consistency-guided incremental tuning
+We focus on fine-tuning only the last $C_{l}$ layers, as the later layers of a pre-trained model are responsible for learning domain-specific fine-grained features, whereas the earlier layers are more general and transferable to a wide range of tasks (Neyshabur et al., 2020).
+Importantly, this fine-tuning is performed under the consistency constraint of asynchronous contrastive learning, ensuring the model learns target-domain features effectively without overfitting or forgetting pre-trained knowledge.
+This balance between adaptability and generalization enables effective adaptation to new classes while preserving the pre-trained encoder’s foundational capabilities.
+The two-setup controlled fine-tuning strategy is applied during the first session, while in later sessions, fine-tuning continues with LoRA across all layers, along with fine-tuning the last $C_{l}$ layers.
+3.2.3 Consistency-guided incremental tuning.
 While the first two modules facilitate tuning the foundation model ($f_{\theta^{{}^{\prime}}}$) during the first session, later sessions also require the retention of previously learned classes.
 To facilitate this, we propose consistency-guided incremental tuning, which prevents forgetting by regularizing the output distribution of the student $f_{\theta^{{}^{\prime}}}$ when training on the incremental sessions.
-More specifically, we enforce consistency between the predictions of the student encoder and a frozen encoder after the first session, effectively discouraging substantial changes in the learned representations of the student.
+More specifically, we enforce consistency between the predictions of the student encoder and the frozen student encoder after the first session, effectively discouraging substantial changes in the learned representations of the student.
+We do not enforce consistency with the frozen encoder from the most recent session (e.g., using the frozen encoder from session $t-1$ for session $t$), since the model may gradually drift away from the pre-trained encoder’s knowledge with cumulative overfitting over the sessions, which can lead to increased forgetting of the foundational generalization capabilities of the model.
+By anchoring the consistency to the first session’s encoder, CoACT effectively balances the learning of new classes with the retention of pre-trained knowledge, mitigating catastrophic forgetting.
+In a standard class-incremental learning setup, the encoder updates across multiple sessions as $f_{\theta_{t}}=f_{\theta^{t-1}}+\Delta_{\theta_{t}}$ where $\Delta_{\theta_{t}}$ represents the parameter updates due to learning new classes at session $t$.
+If is consistency enforced with the encoder from the previous session, $t-1$.
+As a result, the model incrementally adapts but can also gradually drift away from its initial state, leading to $f_{\theta_{T}}=f_{\theta^{0}}+\sum_{t=1}^{T}\Delta_{\theta_{t}}$.
+This, in turn, can result in a gradual loss of generalization, where the model drifts away from the foundational encoder.
+Instead, enforcing consistency with the frozen encoder from the first session $f_{\theta^{0}}$ prevents progressive divergence.
+This ensures that at each session, $f_{\theta_{t}}=f_{\theta^{0}}+\Delta^{\prime}_{\theta_{t}}$, where $\|\Delta^{\prime}_{\theta_{t}}\|\ll\|\sum_{i=1}^{t}\Delta_{\theta_{i}}\|$.
 Let $f_{\theta_{\beta}}$ be the frozen encoder after the first session, and the frozen embedding of this encoder be $p_{i}=f_{\theta_{\beta}}(\mathcal{A}(x_{i}))$.
@@
 4 Experiments
-4.1 Baselines
-In the absence of available FSCIL with foundation model, we first establish a few baselines to better evaluate our proposed framework.
-To this end, we adopt three prominent methods on which to build these baselines.
-First, we use prototype learning, which has been shown to perform well for learning incremental classes in traditional FSCIL [42].
-To apply this approach to a foundation model $f_{\theta}(x)$, we keep the encoder frozen during the incremental learning sessions, while the classifier ($W$) is learned with the prototypes [42] of the new classes: $W=\{w_{0}^{0},w_{1}^{0},\cdot\cdot,w_{|c^{0}|}^{0}\}\cup\{w_{0}^{T},w_{1}^{T},\cdot\cdot,w_{|c^{T}|}^{T}\}$.
+Table 1: Comparison to prior works across the base and incremental sessions on CUB-200 with 100 base classes and 10-way 5-shot evaluation setup.
+Method PT. Backbone Acc. in each session (%) ↑
+0 1 2 3 4 5 6 7 8 9 10
+iCaRL (Rebuffi et al., 2017) 68.68 52.65 48.61 44.16 36.62 29.52 27.83 26.26 24.01 23.89 21.16
+EEIL (Castro et al., 2018) 68.68 53.63 47.91 44.20 36.30 27.46 25.93 24.70 23.95 24.13 22.11
+TOPIC (Tao et al., 2020) 68.68 62.49 54.81 49.99 45.25 41.40 38.35 35.36 32.22 28.31 26.28
+Rebalancing (Hou et al., 2019) 68.68 57.12 44.21 28.78 26.71 25.66 24.62 21.52 20.12 20.06 19.87
+SPPR (Zhu et al., 2021) 68.68 61.85 57.43 52.68 50.19 46.88 44.65 43.07 40.17 39.63 37.33
+MetaFSCIL(Chi et al., 2022) 75.90 72.41 68.78 64.78 62.96 59.99 58.30 56.85 54.78 53.82 52.64
+F2M (Shi et al., 2021) 81.07 78.16 75.57 72.89 70.86 68.17 67.01 65.26 63.36 61.76 60.26
+CEC (Zhang et al., 2021) 75.85 71.94 68.50 63.50 62.43 58.27 57.73 55.81 54.83 53.52 52.28
+FACT (Zhou et al., 2022a) 75.90 73.23 70.84 66.13 65.56 62.15 61.74 59.83 58.41 57.89 56.94
+LIMIT (Zhou et al., 2022b) 75.89 73.55 71.99 68.14 67.42 63.61 62.40 61.35 59.91 58.66 57.41
+SoftNet (Kang et al., 2023) 78.07 74.58 71.37 67.54 65.37 62.60 61.07 59.37 57.53 57.21 56.75
+SAVC (Song et al., 2023) 81.85 77.92 74.95 70.21 69.96 67.02 66.16 65.30 63.84 63.15 62.50
+BOT (Roy et al., 2024) 82.31 78.03 75.45 70.99 71.06 67.85 67.44 66.05 64.95 64.31 63.75
+SV-T (Qiu et al., 2023) SwinT 84.19 82.63 81.21 78.97 79.38 77.64 77.55 75.71 75.91 75.77 76.17
+CPE-CLIP (D’Alessandro et al., 2023) CLIP-B/16 81.58 78.52 76.68 71.86 71.52 70.23 67.66 66.52 65.09 64.47 64.60
+PriViLege (Park et al., 2024) CLIP-B/16 82.21 81.25 80.45 77.76 77.78 75.95 75.69 76.00 75.19 75.19 75.08
+Ours ViT-B/16 88.68 86.26 85.83 83.38 83.52 81.71 81.77 81.77 81.02 80.76 81.19
+Table 2: Comparison to prior works across the base and incremental sessions on miniImageNet with 60 base classes and 5-way 5-shot incremental setting.
+Method PT. Backbone Acc. in each session (%) ↑
+0 1 2 3 4 5 6 7 8
+iCaRL (Rebuffi et al., 2017) 71.77 61.85 58.12 54.60 51.49 48.47 45.90 44.19 42.71
+Rebalancing (Hou et al., 2019) 72.30 66.37 61.00 56.93 53.31 49.93 46.47 44.13 42.19
+TOPIC (Tao et al., 2020) 61.31 50.09 45.17 41.16 37.48 35.52 32.19 29.46 24.42
+EEIL (Castro et al., 2018) 61.31 46.58 44.00 37.29 33.14 27.12 24.10 21.57 19.58
+FSLL (Mazumder et al., 2021) 66.48 61.75 58.16 54.16 51.10 48.53 46.54 44.20 42.28
+FSLL+SS (Mazumder et al., 2021) 68.85 63.14 59.24 55.23 52.24 49.65 47.74 45.23 43.92
+F2M (Shi et al., 2021) 72.05 67.47 63.16 59.70 56.71 53.77 51.11 49.21 47.84
+CEC (Zhang et al., 2021) 72.00 66.83 62.97 59.43 56.70 53.73 51.19 49.24 47.63
+MetaFSCIL (Chi et al., 2022) 72.04 67.94 63.77 60.29 57.58 55.16 52.90 50.79 49.19
+C-FSCIL (Hersche et al., 2022) 76.40 71.14 66.46 63.29 60.42 57.46 54.78 53.11 51.41
+FACT (Zhou et al., 2022a) 72.56 69.63 66.38 62.77 60.60 57.33 54.34 52.16 50.49
+CLOM (Zou et al., 2022) 73.08 68.09 64.16 60.41 57.41 54.29 51.54 49.37 48.00
+LIMIT (Zhou et al., 2022b) 72.32 68.47 64.30 60.78 57.95 55.07 52.70 50.72 49.19
+SoftNet (Kang et al., 2023) 79.77 75.08 70.59 66.93 64.00 61.00 57.81 55.81 54.68
+SAVC (Song et al., 2023) 81.12 76.14 72.43 68.92 66.48 62.95 59.92 58.39 57.11
+BOT (Roy et al., 2024) 84.30 79.59 75.49 71.4 68.45 65.12 62.20 60.52 59.57
+SV-T (Qiu et al., 2023) SwinT 88.75 87.92 86.07 84.84 84.30 83.24 82.22 82.28 82.38
+CPE-CLIP (D’Alessandro et al., 2023) CLIP-B/16 90.23 89.56 87.42 86.80 86.51 85.08 83.43 83.38 82.77
+PriViLege (Park et al., 2024) CLIP-B/16 96.68 96.49 95.65 95.54 95.54 94.91 94.33 94.19 94.10
+Ours ViT-B/16 97.63 97.55 97.09 97.02 97.0 96.58 96.3 96.29 96.24
+4.1 Datasets and implementation details.
+Following existing literature on FSCIL, we evaluate CoACT on CIFAR-100 (Krizhevsky et al., 2009), CUB-200 (Wah et al., 2011), and miniImageNet (Russakovsky et al., 2015) datasets.
+We evaluate our new benchmark, FSCIT on a diverse set of 16 datasets, including generic object detection (Caltech101 (Fei-Fei et al., 2004), CIFAR-100 (Krizhevsky et al., 2009), CUB-200 (Wah et al., 2011), miniImageNet (Russakovsky et al., 2015), VOC 2007 (Everingham, 2008)), fine-grained recognition (OxfordPets (Parkhi et al., 2012), StanfordCars (Krause et al., 2013), Flower102 (Nilsback & Zisserman, 2008), Food101 (Bossard et al., 2014), FGVCAircraft (Maji et al., 2013)), scene recognition (SUN397 (Xiao et al., 2010), Country211 (Radford et al., 2021)), satellite-image (EuroSAT (Helber et al., 2019), Resisc-45 (Cheng et al., 2017)), texture recognition (DTD (Cimpoi et al., 2014)), and traffic sign recognition (GTSRB (Houben et al., 2013)).
+By default, we divide the classes into 10 (or 9) sessions with an equal number of classes and perform 10-shot continual learning.
+The performance after each session is calculated as the average accuracy of all classes seen so far.
+Unless specified otherwise, accuracy refers to the accuracy after the last session, which is the average accuracy over all classes on the test set.
+We also evaluate CoACT on the traditional FSCIL setup, where the initial training is performed with a large in-distribution base session, followed by few-shot tuning over the incremental sessions.
+We use ViT-B/16 as the backbone for most of the experiments, while we present the detailed per-dataset results with ViT-B/32 and ViT-L/16 as well.
+The encoders are pre-trained on ImgeNet-21K (Russakovsky et al., 2015).
+We implement our framework in PyTorch and train the model using an SGD optimizer with a momentum of 0.9.
+The base learning rate is set to 0.1, with a batch size of 64, and the model is trained for 50 epochs for the first session and 5 epochs for the remaining sessions.
+A cosine LR decay scheduler is used to reduce the learning rate over the training epochs.
+The teacher encoder is updated with a momentum value of 0.999.
+For experiments with the FSCIL setup, we train the model for 25 epochs with a learning rate of 0.001.
+We train the model with input resolution of $224\times 224$.
+All other implementation details are the same as described above.
+All experiments are conducted with 3 random seeds, and the reported results are averaged over the three runs.
+All experiments are conducted on an Nvidia V100 GPU, where the training takes about 6 hours.
+Table 3: Comparison to prior works across the base and incremental sessions on CIFAR-100 with 60 base classes and 5-way 5-shot incremental setting.
+Method PT. Backbone Acc. in each session (%) ↑
+0 1 2 3 4 5 6 7 8
+Rebalancing (Hou et al., 2019) 61.31 47.80 39.31 31.91 25.68 21.35 18.67 17.24 14.17
+iCaRL (Rebuffi et al., 2017) 61.31 46.32 42.94 37.63 30.49 24.00 20.89 18.80 17.21
+TOPIC (Tao et al., 2020) 61.31 50.09 45.17 41.16 37.48 35.52 32.19 29.46 24.42
+IDLVQ-C (Chen & Lee, 2020) 64.77 59.87 55.93 52.62 49.88 47.55 44.83 43.14 41.84
+FSLL (Mazumder et al., 2021) 66.48 61.75 58.16 54.16 51.10 48.53 46.54 44.20 42.28
+FSLL+SS (Mazumder et al., 2021) 68.85 63.14 59.24 55.23 52.24 49.65 47.74 45.23 43.92
+F2M (Shi et al., 2021) 67.28 63.80 60.38 57.06 54.08 51.39 48.82 46.58 44.65
+CEC (Zhang et al., 2021) 73.07 68.88 65.26 61.19 58.09 55.57 53.22 51.34 49.14
+MetaFSCIL (Chi et al., 2022) 74.50 70.10 66.84 62.77 59.48 56.52 54.36 52.56 49.97
+CLOM (Zou et al., 2022) 74.2 69.83 66.17 62.39 59.26 56.48 54.36 52.16 50.25
+C-FSCIL (Hersche et al., 2022) 77.47 72.40 67.47 63.25 59.84 56.95 54.42 52.47 50.47
+LIMIT (Zhou et al., 2022b) 73.81 72.09 67.87 63.89 60.70 57.77 55.67 53.52 51.23
+FACT (Zhou et al., 2022a) 74.60 72.09 67.56 63.52 61.38 58.36 56.28 54.24 52.10
+SAVC (Song et al., 2023) 79.85 73.70 69.37 65.28 61.91 59.27 57.24 54.97 53.12
+SoftNet (Kang et al., 2023) 79.88 75.54 71.64 67.47 64.45 61.09 59.07 57.29 55.33
+BOT (Roy et al., 2024) 80.25 77.20 75.09 70.82 67.83 64.86 62.73 60.52 58.75
+SV-T (Qiu et al., 2023) SwinT 86.77 82.82 80.36 77.20 76.06 74.00 72.92 71.68 69.75
+CPE-CLIP (D’Alessandro et al., 2023) CLIP-B/16 87.83 85.86 84.93 82.85 82.64 82.42 82.27 81.44 80.52
+PriViLege (Park et al., 2024) CLIP-B/16 90.88 89.39 88.97 87.55 87.83 87.35 87.53 87.15 86.06
+Ours ViT-B/16 90.46 88.46 88.11 86.94 86.98 86.52 86.39 86.0 84.63
+4.2 FSCIL results
+In this section, we discuss the results of CoACT in the FSCIL setup.
+Following the existing literature, we present these results on the CIFAR-100, CUB-200, and miniImageNet datasets.
+In Table 1, we compare the performance of CoACT with prior works on the CUB-200 dataset, where we divide the existing methods into two groups: the first group, which trains randomly initialized models and the second group which use a pre-trained encoder.
+The results are reported for 100 base classes and a 10-way, 5-shot incremental learning setup.
+As we observe, CoACT outperforms the previous SOTA (SV-T) by 5.02%.
+Results on miniImageNet are reported in Table 2, with 60 base classes and a 5-way, 5-shot incremental learning setup.
+As we find from this table, CoACT outperforms the prior SOTA (PriViLege) by 2.14%.
+Also, the performance difference increases (over the sessions) as we learn new classes, indicating that CoACT shows less forgetting than PriViLege.
+Finally, the results for CIFAR-100 are presented in Table 3, where PriViLege (Park et al., 2024) holds the highest result.
+With significantly fewer parameters (86M in ViT-B/16 compared to 149M in CLIP-B/16), our proposed CoACT achieves competitive results, performing only 1.43% below PriViLege.
+Dataset Classes Sessions First Ses. class Cls/session
+Caltech101 102 10 12 10
+CIFAR-100 100 10 10 10
+Country211 211 10 22 21
+CUB-200 200 10 20 20
+DTD 47 9 5 5
+EuroSAT 10 10 1 1
+FGVCAircraft 100 10 10 10
+Food101 101 10 11 10
+GTSRB 43 10 7 4
+MiniImageNet 100 10 10 10
+Flower102 102 10 12 10
+OxfordPets 37 9 5 4
+Resisc-45 45 9 5 5
+StanfordCars 196 9 28 21
+SUN397 397 10 45 44
+VOC 2007 20 10 2 2
+Table 4: Details on class splits over the continual sessions for different datasets.
+Method
+Caltech101
+CIFAR100
+Country211
+CUB200
+DTD
+EuroSat
+FGVCAircraft
+Food101
+GTSRB
+MiniImageNet
+Flowers102
+OxfordPets
+Resisc-45
+StanfordCars
+SUN397
+VOC 2007
+Average
+CPE-CLIP 82.79 74.11 7.03 73.47 53.92 61.77 14.79 69.71 14.64 94.57 97.02 87.05 48.43 18.51 64.48 63.53 57.85
+PriViLege 83.49 75.06 7.28 74.45 54.44 62.57 15.23 70.67 15.87 95.04 97.46 87.34 49.17 18.67 65.09 64.11 58.49
+Pro. lear. 83.47 75.58 7.42 74.10 54.72 62.20 15.51 71.09 15.53 95.31 98.09 88.10 49.83 18.90 65.68 64.72 58.77
+Lin. tun. 83.60 75.58 7.45 74.41 54.72 62.20 15.32 71.18 16.92 95.32 98.09 88.10 49.83 18.96 65.87 64.72 58.89
+LoRA 83.87 75.81 7.45 74.65 54.70 62.20 15.32 71.24 17.23 95.32 98.09 88.10 49.82 18.94 65.90 64.72 58.96
+CoACT 86.86 78.31 7.42 77.38 55.11 62.25 18.98 71.59 26.05 95.34 98.18 87.79 62.62 24.40 66.11 64.45 61.43
+Table 5: Performance of CoACT on FSCIT across 16 datasets and its comparison to existing FSCIL methods and our baseline methods.
+Pro.
+lear.
+and Lin.
+tun.
+respectively refer to prototype learning and linear tuning.
+4.3 FSCIT results
+As the first work on FSCIT, we first establish a few baselines to better evaluate our proposed framework.
+To this end, we explore two of the existing SOTA from the FSCIL literature PriViLege and CPE-CLIP and also explore 3 well-known approaches from the few-shot tuning literature.
+First, we adopt prototype learning, which has been shown to perform well for learning incremental classes in traditional FSCIL (Song et al., 2023).
+To apply this approach to FSCIT, we keep the pre-trained encoder frozen during the incremental learning sessions, while the classifier ($W$) is learned with the prototypes (Song et al., 2023) of the new classes: $W=\{w_{0}^{0},w_{1}^{0},\cdot\cdot,w_{|c^{0}|}^{0}\}\cup\{w_{0}^{T},w_{1}^{T},\cdot\cdot,w_{|c^{T}|}^{T}\}$.
 Here, the prototype $w_{c}^{t}$ of class $c$ from session $t$ is defined as the average of the embeddings of all samples of the class: ${w}_{c}^{t}=\frac{1}{|P_{i}|}\sum_{j\in P_{i}}{f_{\theta}({{x}}_{j})}$, where $P_{i}\stackrel{{\scriptstyle\text{def}}}{{=}}\{j:y_{j}=i\}$.
-Next, we consider linear tuning as a baseline and train the model in an incremental-frozen approach, where the encoder is trained only on the first session, followed by prototype learning for the remaining sessions [50].
-Finally, we use parameter-efficient fine-tuning as the third baseline, where we add LoRA to different layers of the pre-trained encoder and train newly added LoRA while keeping the encoder frozen.
-4.2 Experiment setup
-Dataset.
-We comprehensively evaluate the performance of our proposed method on a diverse set of 16 datasets, including generic object detection (Caltech101 [12], CIFAR-100 [26], CUB-200 [45], MiniImageNet [40], VOC 2007 [11]), fine-grained recognition (OxfordPets [34], StanfordCars [25], Flower102 [33], Food101 [1], FGVCAircraft [29]), scene recognition (SUN397 [46], Country211 [36]), satellite-image (EuroSAT [14], Resisc-45 [5]), texture recognition (DTD [8]), and traffic sign recognition (GTSRB [17]).
-Method Acc
-Prototype learning 58.77$\pm$0.0
-Linear tuning 58.89$\pm$0.23
-LoRA 58.96$\pm$0.11
-CoACT 61.43$\pm$0.12
-Table 1: Overall performance over 16 datasets.
-Implementation details.
-By default, we divide the classes into 10 (or 9; see Appendix for details) sessions with an equal number of classes and perform 10-shot continual training following FSCIL literature [50; 51].
-The performance after each session is calculated as the average accuracy of all classes seen so far.
-Unless specified otherwise, accuracy refers to the accuracy after the last session, which is the average accuracy over all classes.
-We also evaluate CoACT on the traditional FSCIL setup, where the initial training is performed with a large labelled set from the same distribution as the incremental sessions, followed by tuning on the few-shot incremental sessions.
-We use a ViT-B/16 as the backbone for most of the experiments, while we present the detailed per-dataset results with ViT-B/32 and ViT-L/16 as well.
-The encoders are pre-trained on ImgeNet-21K [40].
-The framework is implemented on PyTorch, and experiments are conducted on an Nvidia V100 GPU, where the training takes about 6 hours.
-More details on the hyper-parameters are provided in Appendix A.2.
-4.3 Results
-Figure 5: (left) Forgetting of learned classes. (right) Accuracy breakdown over the first session, remaining sessions, and all sessions.
-Table 1 presents the average accuracy and standard deviation (over 3 runs) of our method over 16 datasets, in comparison to the aforementioned baselines.
-As we find from the table, CoACT achieves an average accuracy of 61.43% over the 16 datasets, outperforming others by an average of 2.47%.
-Among the baselines, LoRA shows the second-best accuracy with 58.96%.
-The other two baselines show relatively similar average accuracies to LoRA.
-Note that among the baselines, prototype learning with a frozen encoder is a deterministic process, thus a standard deviation of 0.
-Note that as we will later demonstrate in Table 3, the SOTA in the standard FSCIL setup achieve $\sim$1gains versus prior work.
-Table 2: Performance of CoACT on 16 datasets and its comparison to the baseline methods.
-Method
+Next, we consider an incremental-frozen approach with linear tuning as a baseline, where we fine-tune the model only on the first session, followed by prototype learning for the remaining sessions (Zhou et al., 2022a).
+Finally, we use parameter-efficient fine-tuning as the third baseline, where we add LoRA to different layers of the pre-trained encoder, train them during the first setting, and perform prototype learning in incremental sessions.
+Additionally, to ensure extensive evaluation of the new benchmark, we conduct the experiments on 16 datasets.
+The classes are divided into sessions as summarized in Table 4.
+In Table 5, we present the main results of our study on the FSCIT setup using the ViT-B/16 encoder.
+The table highlights that existing methods for FSCIL, such as CPE-CLIP and PriViLege, do not perform well on FSCIT.
+This is primarily because these methods heavily rely on supervised training of the base session, which is absent in the FSCIT setup, leading to reduced performance.
+Among the other three baselines, namely prototype tuning, linear tuning, and LoRA, performance is slightly better compared to FSCIL methods.
+However, CoACT surpasses all prior approaches by significant margins on all the datasets.
+Specifically, CoACT achieves an average improvement of 2.47% over the baselines and exhibits notable gains of up to 12.79% on individual datasets, such as Resisc-45.
+Additionally, we observe larger improvements on more challenging datasets.
+For instance, on the five datasets with the lowest accuracy (Country211, FGVCAircraft, GTSRB, Resisc-45, and StanfordCars), CoACT achieves an average improvement of 6.14% over the best-performing baseline.
 Encoder
@@
 Average
-Pro. lear. ViT-B/16 83.47 75.58 7.42 74.10 54.72 62.20 15.51 71.09 15.53 95.31 98.09 88.10 49.83 18.90 65.68 64.72 58.77
-ViT-B/32 83.90 74.29 6.79 69.32 52.83 68.91 13.50 63.57 19.84 92.76 97.09 84.57 45.30 16.85 61.29 66.64 57.34
-ViT-L/16 85.10 75.04 7.23 76.31 54.33 65.48 19.03 74.66 12.24 98.05 98.82 87.64 54.85 24.69 66.58 60.16 60.01
-Lin. tun. ViT-B/16 83.60 75.58 7.45 74.41 54.72 62.20 15.32 71.18 16.92 95.32 98.09 88.10 49.83 18.96 65.87 64.72 58.89
-ViT-B/32 83.87 74.35 6.75 69.84 53.22 68.89 11.29 63.51 26.28 92.75 97.64 85.06 51.17 16.57 61.36 66.58 58.07
-ViT-L/16 85.25 75.33 7.26 76.11 54.24 65.32 19.44 74.51 13.01 97.61 98.34 87.34 54.99 25.21 66.99 61.01 60.12
-LoRA ViT-B/16 83.87 75.81 7.45 74.65 54.70 62.20 15.32 71.24 17.23 95.32 98.09 88.10 49.82 18.94 65.90 64.72 58.96
-ViT-B/32 83.54 74.72 6.81 69.85 53.22 68.89 10.88 63.52 29.84 92.75 97.55 85.09 49.14 16.09 61.32 66.56 58.11
-ViT-L/16 85.15 75.34 7.29 76.26 54.52 65.68 19.13 74.75 12.28 97.88 98.71 87.89 54.85 24.99 66.78 60.56 60.13
-CoACT ViT-B/16 86.86 78.31 7.42 77.38 55.11 62.25 18.98 71.59 26.05 95.34 98.18 87.79 62.62 24.40 66.11 64.45 61.43
+ViT-B/16 86.86 78.31 7.42 77.38 55.11 62.25 18.98 71.59 26.05 95.34 98.18 87.79 62.62 24.40 66.11 64.45 61.43
 ViT-B/32 85.59 77.26 6.85 72.21 54.78 67.71 14.56 68.41 29.99 94.81 98.01 86.21 55.34 20.18 64.49 66.71 60.19
-ViT-L/16 87.25 79.27 7.45 78.29 56.19 66.36 21.24 74.99 17.22 97.18 98.18 87.95 62.25 25.43 67.18 62.23 61.79
-[…]
-[…]
-Figure 6: Performance vs. different shots.
-Figure 7: Comparison to SOTA on traditional FSCIL.
-Figure 8: Forgetting over incr. sessions.
+ViT-L/14 87.25 79.27 7.45 78.29 56.19 66.36 21.24 74.99 17.22 97.18 98.18 87.95 62.25 25.43 67.18 62.23 61.79
+Swin-B 88.01 80.17 8.15 78.91 57.25 68.98 23.44 76.11 28.18 97.01 98.03 88.56 64.24 25.90 69.34 66.95 63.70
+DINO ViT-B/14 87.87 79.92 7.89 79.01 57.81 68.57 22.58 76.47 30.38 96.94 98.05 88.29 63.59 25.87 69.04 65.47 63.61
+CLIP ViT-B/16 (V) 84.23 78.14 6.49 72.89 55.23 68.48 24.93 72.49 27.38 95.18 97.23 86.66 57.23 23.47 64.29 62.38 61.04
+Table 6: Performance of CoACT on FSCIT across 16 datasets on different encoders.
+Figure 4: (left) Forgetting of learned classes in FSCIT. (right) Accuracy breakdown into the first, remaining and all sessions in FSCIT setup.
 Next, we report the forgetting of learned classes for each method, measured as the drop in accuracy w.r.t.
 the first session, in Figure 4 (left).
-As we observe in this figure, CoACT has the least amount of forgetting compared to the baselines with approximately 1.5% less forgetting than both LoRA and linear tuning, and 3.2% than prototype learning.
+As we observe in this figure, CoACT has the least amount of forgetting compared to the baselines, with approximately 1.5% less forgetting than both LoRA and linear tuning and 3.2% than prototype learning.
 We also present a breakdown of accuracies into the first incremental session, the remaining incremental sessions, and all sessions in Figure 4 (right).
@@
 In contrast, CoACT shows higher improvement in both the first and remaining sessions.
-To further evaluate the efficacy of CoACT, we investigate the performance with a different number of samples per class, including 1, 2, 4, 8, and 16.
-As seen in Figure 8, CoACT outperforms other methods in all settings and shows robustness in very-low-shot settings.
+Figure 5: Performance for different shots in FSCIT.
+Next, we investigate the generalization of CoACT across different encoders of different sizes.
+Specifically, we explore ViT-B/16, ViT-B/32, ViT-L/14, Swin-B (Liu et al., 2021), DINO ViT-B/14 (Oquab et al., ), and vision encoder from CLIP ViT-B/16 (Radford et al., 2021) as the backbone.
+The results of this study are presented in Table 6.
+The results from this study show that our proposed solution generalizes across the encoder sizes.
+Specifically, with a larger encoder (ViT-L/14), CoACT improves the average accuracy to 61.79%.
+On individual datasets, we find up to 3.4% improvements over the default ViT-B/16 encoder.
+Similarly, CoACT shows strong performances on the Swin-B, DINO ViT-B/14, and CLIP ViT-B/16 encoders.
+To further evaluate the efficacy of CoACT, we investigate the performance with a different number of samples per class.
+Specifically, we explore fine-tuning the foundation model in the FSCIT setup with only 1, 2, 4, 8, and 16 samples per class.
+The results of this experiment are presented in Figure 5.
+As we find from this experiment, CoACT outperforms other methods in all settings and shows robustness in very low-shot settings.
 The baselines vary in effectiveness with the different number of samples per class.
 For instance, prototype learning performs better than linear tuning and LoRA in 1-shot and 2-shot settings, while the other two perform better as the number of samples increases.
-These results indicate that linear tuning and LoRA do not learn very effective representations in very-low-shot settings, compared to our method and prototype learning.
-Finally, in Table 2, we present per-dataset results using three different encoder backbones.
-With ViT-B/16 as the backbone, while CoACT shows average improvements of 2.66%, 2.54%, and 2.47% improvement over prototype learning, linear tuning, and LoRA, respectively.
-Notably, it shows up to a 12.79% improvement on some individual datasets such as Resisc-45.
-In general, we find relatively higher improvements for more challenging datasets.
-For instance, on the 5 datasets with the least accuracy (Country211, FGVCAircraft, GTSRB, Resisc-45, and StanfordCars), CoACT shows an average improvement of 6.14% over the best baseline method (LoRA).
-We observe similar improvements in the performance with the other two encoders (ViT-B/32 and ViT-L/16), where CoACT outperforms the best baseline (LoRA) by average accuracies of 2.08% and 1.67%.
-To further evaluate the effectiveness of our method, we also investigate its performance in the traditional training setup of FSCIL.
-Following the existing literature, we present these results on the CIFAR-100, CUB-200, and miniImageNet datasets.
-In Table 3, we compare the performance of CoACT with prior works on the CIAFR-100 dataset, where we divide the existing methods into two groups: the first group, which trains randomly initialized models and the second group which use a pre-trained encoder.
-An extension of this table with more prior works is presented in the Appendix A.3.
-As we observe in the table, CoACT outperforms the previous SOTA without a pre-trained encoder (BOT [39]) by 25.88% and the SOTA with a pre-trained encoder (CPE-CLIP [9]) by 4.11%.
-For a fair comparison with existing methods, we evaluate SAVC [42] and BOT [39] with the same pre-trained encoder (ViT-B/16) as ours, where both methods show considerably lower performance compared to CoACT.
-Results on the CUB-200 and miniImageNet datasets are presented in Figure 8 (See Appendix A.3).
-For the CUB-200 dataset, the proposed method outperforms BOT and CPE-CLIP by 17.44% and 5.02%, respectively, while the improvements over these two SOTA methods on the miniImageNet dataset are 36.67% and 13.47%.
-Table 3: Comparison to prior works on traditional FSCIL on CIFAR-100 with 60 base classes and 5-way 5-shot incremental setting.
-We reproduce existing SOTA methods (marked with *) with the same backbone as CoACT.
-Results on the CUB-200 and miniImageNet are presented in the Appendix.
-Method Pre-trained Enc. Acc. in each session (%) ↑
-0 1 2 3 4 5 6 7 8
-CEC [48] 73.07 68.88 65.26 61.19 58.09 55.57 53.22 51.34 49.14
-MetaFSCIL [7] 74.50 70.10 66.84 62.77 59.48 56.52 54.36 52.56 49.97
-CLOM [53] 74.2 69.83 66.17 62.39 59.26 56.48 54.36 52.16 50.25
-C-FSCIL [15] 77.47 72.40 67.47 63.25 59.84 56.95 54.42 52.47 50.47
-LIMIT [51] 73.81 72.09 67.87 63.89 60.70 57.77 55.67 53.52 51.23
-FACT [50] 74.60 72.09 67.56 63.52 61.38 58.36 56.28 54.24 52.10
-SAVC [42] 79.85 73.70 69.37 65.28 61.91 59.27 57.24 54.97 53.12
-SoftNet [21] 79.88 75.54 71.64 67.47 64.45 61.09 59.07 57.29 55.33
-BOT [39] 80.25 77.20 75.09 70.82 67.83 64.86 62.73 60.52 58.75
-SAVC* [42] ViT-B/16 82.98 75.35 74.01 71.51 70.64 69.78 68.98 67.84 66.24
-BOT* [39] ViT-B/16 83.75 78.14 76.85 73.23 72.95 72.03 71.56 70.66 69.72
-SV-T [35] SwinT 86.77 82.82 80.36 77.20 76.06 74.00 72.92 71.68 69.75
-CPE-CLIP [9] CLIP-B/16 87.83 85.86 84.93 82.85 82.64 82.42 82.27 81.44 80.52
-CoACT (Ours) ViT-B/16 90.46 88.46 88.11 86.94 86.98 86.52 86.39 86.0 84.63
-Table 4: Performance on taditional FSCIL with a few-shot base session.
-Method Base shots Acc. in each session (%) ↑
-0 1 2 3 4 5 6 7 8
-BOT [39] 500 80.25 77.20 75.09 70.82 67.83 64.86 62.73 60.52 58.75
-CPE-CLIP[9] 500 87.83 85.86 84.93 82.85 82.64 82.42 82.27 81.44 80.52
-CoACT (Ours) 1 59.04 57.79 58.91 58.65 59.46 59.64 60.01 60.42 59.91
-4 72.9 71.11 71.39 70.75 71.24 70.95 71.2 71.16 70.57
-8 76.53 74.78 74.71 73.95 74.34 73.89 73.99 73.81 73.05
-32 83.87 82.0 81.61 80.24 80.27 79.64 79.56 79.25 78.35
-128 83.93 83.1 82.69 82.21 81.94 81.54 81.16 80.90 80.69
-Next, we investigate how CoACT compares to previous SOTA methods, namely BOT and CPE-CLIP, in terms of forgetting the already learned classes.
-As we find from Figure 8, CoACT shows a lower forgetting than previous SOTA methods over the incremental sessions.
-Specifically, CoACT shows a forgetting of 5.8% compared to the 7.3% of CPE-CLIP and 21.5% of BOT.
-This further shows the robustness of CoACT against forgetting during incremental learning sessions.
-ACT C.F. C.I.T Acc
-61.43
-61.31
-61.28
-60.26
-58.77
-Table 5: Ablation study.
-We also study how CoACT performs with a reduced number of samples per base class in the traditional FSCIL setup.
-For this experiment, we keep the incremental session setup the same (5-way 5-shot for CIFAR-100) but use only a few samples per class for the base session.
-The results of this study are presented in Table 4.
-As we find from this experiment, with only 1 sample per base class, CoACT achieves an accuracy of 59.91% in the final session, outperforming BOT with 500 samples per class.
-We also outperform CPE-CLIP with an accuracy of 80.69% in the final session when we use 128 samples per base class.
+These results indicate that linear tuning and LoRA do not learn very effective representations in very low-shot settings compared to our method and prototype learning.
+Table 7: Ablation study on all datasets.
+Here, Asy.
+Cont., Cal.
+Tun., Con.
+tun.
+refers to asynchronous contrastive tuning, controlled fine-tuning and consistency-guided incremental tuning.
+The reported values are average and standard deviations across three runs.
+Asy. Cont.
+Cal. Tun.
+Con. Tun.
+Caltech101
+CIFAR100
+Country211
+CUB200
+DTD
+EuroSat
+FGVCAircraft
+Food101
+$86.86{\scriptstyle\pm 0.03}$ $78.31{\scriptstyle\pm 0.02}$ $7.42{\scriptstyle\pm 0.04}$ $77.38{\scriptstyle\pm 0.01}$ $55.11{\scriptstyle\pm 0.05}$ $62.25{\scriptstyle\pm 0.03}$ $18.98{\scriptstyle\pm 0.02}$ $71.59{\scriptstyle\pm 0.08}$
+$85.89{\scriptstyle\pm 0.05}$ $78.45{\scriptstyle\pm 0.03}$ $7.45{\scriptstyle\pm 0.02}$ $77.27{\scriptstyle\pm 0.04}$ $54.59{\scriptstyle\pm 0.01}$ $62.91{\scriptstyle\pm 0.05}$ $18.20{\scriptstyle\pm 0.03}$ $71.83{\scriptstyle\pm 0.07}$
+$86.01{\scriptstyle\pm 0.02}$ $78.96{\scriptstyle\pm 0.01}$ $7.46{\scriptstyle\pm 0.03}$ $77.20{\scriptstyle\pm 0.05}$ $54.82{\scriptstyle\pm 0.02}$ $62.31{\scriptstyle\pm 0.01}$ $18.91{\scriptstyle\pm 0.09}$ $71.47{\scriptstyle\pm 0.03}$
+$84.62{\scriptstyle\pm 0.12}$ $78.37{\scriptstyle\pm 0.04}$ $7.49{\scriptstyle\pm 0.01}$ $76.44{\scriptstyle\pm 0.03}$ $54.83{\scriptstyle\pm 0.05}$ $62.22{\scriptstyle\pm 0.02}$ $18.23{\scriptstyle\pm 0.01}$ $71.39{\scriptstyle\pm 0.04}$
+$83.47{\scriptstyle\pm 0.10}$ $75.58{\scriptstyle\pm 0.02}$ $7.42{\scriptstyle\pm 0.05}$ $74.10{\scriptstyle\pm 0.01}$ $54.72{\scriptstyle\pm 0.03}$ $62.20{\scriptstyle\pm 0.04}$ $15.51{\scriptstyle\pm 0.02}$ $71.09{\scriptstyle\pm 0.05}$
+Asy. Cont.
+Cal. Tun.
+Con. Tun.
+MiniImageNet
+Flowers102
+OxfordPets
+Resisc-45
+StanfordCars
+SUN397
+VOC 2007
+GTSRB
+Average
+$95.34{\scriptstyle\pm 0.03}$ $98.18{\scriptstyle\pm 0.02}$ $87.79{\scriptstyle\pm 0.01}$ $62.62{\scriptstyle\pm 0.04}$ $24.40{\scriptstyle\pm 0.05}$ $66.11{\scriptstyle\pm 0.03}$ $64.45{\scriptstyle\pm 0.02}$ $26.05{\scriptstyle\pm 0.03}$ 61.43
+$95.04{\scriptstyle\pm 0.01}$ $98.01{\scriptstyle\pm 0.03}$ $88.25{\scriptstyle\pm 0.02}$ $61.77{\scriptstyle\pm 0.05}$ $24.14{\scriptstyle\pm 0.01}$ $66.01{\scriptstyle\pm 0.04}$ $64.89{\scriptstyle\pm 0.03}$ $26.29{\scriptstyle\pm 0.04}$ 61.31
+$95.34{\scriptstyle\pm 0.02}$ $98.15{\scriptstyle\pm 0.01}$ $87.92{\scriptstyle\pm 0.03}$ $61.31{\scriptstyle\pm 0.02}$ $24.07{\scriptstyle\pm 0.04}$ $65.98{\scriptstyle\pm 0.05}$ $64.59{\scriptstyle\pm 0.01}$ $26.05{\scriptstyle\pm 0.02}$ 61.28
+$95.34{\scriptstyle\pm 0.04}$ $98.09{\scriptstyle\pm 0.05}$ $88.10{\scriptstyle\pm 0.01}$ $57.21{\scriptstyle\pm 0.03}$ $23.18{\scriptstyle\pm 0.02}$ $65.94{\scriptstyle\pm 0.01}$ $64.72{\scriptstyle\pm 0.04}$ $18.02{\scriptstyle\pm 0.09}$ 60.26
+$95.31{\scriptstyle\pm 0.05}$ $98.09{\scriptstyle\pm 0.02}$ $88.10{\scriptstyle\pm 0.04}$ $49.83{\scriptstyle\pm 0.01}$ $18.90{\scriptstyle\pm 0.03}$ $65.68{\scriptstyle\pm 0.02}$ $64.72{\scriptstyle\pm 0.05}$ $15.53{\scriptstyle\pm 0.08}$ 58.77
 4.4 Ablation study
-We present an ablation study on the proposed components of CoACT in Table 5, where ACT, C.F., and C.I.T stand for asynchronous contrastive tuning, controlled fine-tuning, and consistency-guided incremental tuning, respectively.
+We present an ablation study on the proposed components of CoACT in Table 7.
 Given that the asynchronous contrastive tuning component of our method could not be removed as it contains the trainable parameters, we start this study by removing controlled fine-tuning and consistency-guided incremental tuning modules individually and simultaneously.
@@
 Finally, with the ablation of all three components and only training a linear classifier, we observe a 2.66% drop in performance.
+Notably, all the ablations experiments show a very small standard deviation.
 Next, we present a comprehensive study on different parameters of CoACT in Table 8.
-In the first study (Table 6(a)), we study different alternates for LoRA, namely Adapter [18] and Prompt [27], where we observe that Adapter and Prompt do not perform as well as LoRA, with a final accuracy of 60.01% and 59.87% versus 60.15%, respectively.
+In the first study (Table 8(a)), we study different alternates for LoRA, namely Adapter (Houlsby et al., 2019) and Prompt (Jia et al., 2022), where we observe that Adapter and Prompt do not perform as well as LoRA, with a final accuracy of 60.01% and 59.87% versus 60.15%, respectively.
+In this table, the prompt implementation follows the visual prompt tuning (VPT) (Jia et al., 2022) approach, where learnable prompts are added to the input embeddings of the vision transformer.
 In the next table (Table 8(b)), we study the performance for different numbers of training epochs, where we find the best results when trained for 50 epochs.
@@
 Moreover, the best result is obtained when starting the fine-tuning step after 10 epochs of training of the newly added LoRA layers.
-Lastly, In Table 6(h), we show the results for training different layers during incremental learning where we find the best results for tuning both the classifier and the LoRA layers during incremental tuning.
+Table 8(h) shows the results for training different layers during incremental learning, where tuning both the encoder (last few layers) and the LoRA shows the best results.
+Table 8(i) shows the study on using the encoder from the first session vs.
+the encoder from the most recent session to enforce consistency.
+As we find from this table, the student encoder from the first session results in a higher performance.
+We also study the performance of two new variants of LoRA, including qLoRA (Dettmers et al., 2024) and LoRA+ (Hayou et al., 2024) in Table 8(j).
+As we find from this table, LoRA+ shows a minor improvement over the Naive LoRA, while qLoRA results in a slight drop in performance.
+However, in our work, we used the default LoRA method given its popularity and wide use in the area as the default approach to fine-tuning.
 Module Accuracy
@@
 Tuning layer Accuracy
-Classifier (clf) 61.22
-clf + LoRA 61.28
-LoRA 60.94
-Table 6: Ablation of different parameters of CoACT on FSCIL averaged across 16 datasets. Full experiments on all datasets are presented in the Appendix.
+Encoder only 61.22
+LoRA only 60.94
+Enc. + LoRA 61.28
+Setup Accuracy
+Consistency w/ first 61.43
+Consistency w/ last 61.22
+Setup Accuracy
+LoRA 58.96
+qLoRA 58.92
+LoRA+ 59.11
+Method Param. Thr. (samp./sec.)
+Train Infer.
+PriViLege 149M 617 1236
+CoACT 86M 930 1850
+Table 8: Ablation of different parameters of CoACT on FSCIL averaged across 16 datasets, including different alternate for LoRA, training epochs, number of LoRA layers, settings for adding LoRA to the encoders, different parameters for controlled tuning (learning rate factor, fine-tuning layers, and fine-tuning epochs), and different tuning layers.
+4.5 Computational complexity
+Finally, we analyze the computational complexity of our method compared to the previous SOTA, PriViLege (Park et al., 2024).
+Our analysis is presented in Table 8(k).
+In terms of the number of parameters, CoACT has significantly fewer parameters than PriViLege (86 million vs.
+149 million) since CoACT is based on a vision-only backbone, whereas PriViLege utilizes a vision-language backbone.
+Additionally, CoACT achieves higher throughput than PriViLege during both training and inference.
+Specifically, PriViLege achieves a training throughput of 617 samples/second, whereas CoACT reaches 930.
+Similarly, the inference throughput is 1,236 for PriViLege and 1,850 for CoACT.
 5 Conclusion
@@
 Extensive studies show the effectiveness of our method, achieving higher accuracy, lower forgetting, and robustness in low-shot settings.
-CoACT also outperforms prior SOTA in the standard FSCIL setup by a large margin.
-We present comprehensive experiments on different components of CoACT and make our code available to foster rapid developments in the area.
+Notably, CoACT outperforms prior works by up to 5.02% in standard FSCIL setup and by up to 12.51% on FSCIT in the individual datasets.
+We present comprehensive experiments on different components of CoACT.
 Limitations.
@@
 On the other hand, the proposed solution can potentially have a large positive impact, as the idea proposed in this work can be explored to update foundation models for new classes without the need to train from scratch, saving compute and other resources.
-Appendix A Appendix
-A.1 Datasets
-Dataset Classes Sessions First Ses. class Cls/session
-Caltech101 102 10 12 10
-CIFAR-100 100 10 10 10
-Country211 211 10 22 21
-CUB-200 200 10 20 20
-DTD 47 9 5 5
-EuroSAT 10 10 1 1
-FGVCAircraft 100 10 10 10
-Food101 101 10 11 10
-GTSRB 43 10 7 4
-MiniImageNet 100 10 10 10
-Flower102 102 10 12 10
-OxfordPets 37 9 5 4
-Resisc-45 45 9 5 5
-StanfordCars 196 9 28 21
-SUN397 397 10 45 44
-VOC 2007 20 10 2 2
-Table 7: Details on class splits over the continual sessions for different datasets
-To comprehensively evaluate CoACT, we run the experiments on a set of 16 diverse datasets.
-Following the traditional setup of FSCIL [51; 50; 44], we divide each dataset into 10 sessions with equal classes where possible; otherwise, we divide them into 9 classes.
-In Table 7, we summarize the details of the total number of classes, number of sessions, and classes per session for all datasets.
-Training is conducted on a few shot samples from the training set of the corresponding dataset, while the accuracies are reported in the whole validation split of the corresponding datasets.
-The few-shot samples are constant across different experiments to ensure a fair comparison and are provided with our codebase.
-A.2 Implementation details
-We implement our framework in PyTorch and train the model using an SGD optimizer with a momentum of 0.9.
-The base learning rate is set to 0.1, with a batch size of 64, and the model is trained for 50 epochs for the first session and 5 epochs for the remaining sessions.
-A cosine LR decay scheduler is used to reduce the learning rate over the training epochs.
-The teacher encoder is updated with a momentum value of 0.999.
-For experiments with the traditional FSCIL setup, we train the model for 25 epochs with a learning rate of 0.001.
-All other implementation details are the same as described above.
-All experiments are conducted with 3 random seeds, and the reported results are averaged over the three runs.
-A.3 Additional results
-In Table 8, we extend the results from Table 1 to summarize the results of CoACT and its comparison to the baselines for all datasets over the continual learning sessions.
-In Table 9, we present a comparison to prior works on traditional FSCIL on CIFAR-100 with 60 base classes.
-This is the extension of Table 3 to include more prior works.
-In Table 10, we present the comparison to prior works on traditional FSCIL on CUB-200 with 100 base classes.
-Comparison to prior works on traditional FSCIL on miniImageNet with 60 base classes is presented in Table 11.
-Complete ablation studies across all 16 datasets are presented in Table 12.
-This table is an extension of Table 5 from the main paper.
-In Figure 9, we present a study on different parameters of asynchronous contrastive tuning, including the number of training epochs and the number of adapter layers.
-Next, in Figure 10, we summarize the study on different parameters of controlled fine-tuning, including learning rate, learning rate factor, encoder fine-tuning start layer, and encoder fine-tuning start epoch.
-Finally, in Figure 11, we show the study on incremental fine-tuning parameters, $\lambda$, and $\gamma$.
-Table 8: Performance of ACT on 16 datasets over the incremental sessions, and its comparison to the baseline methods.
-Dataset Method 1 2 3 4 5 6 7 8 9 10
-CAL101 Prototype learning 82.29 84.55 82.63 76.35 76.43 78.77 82.79 82.82 82.99 83.47
-Linear tuning 83.49 85.0 82.89 76.5 76.56 78.88 82.89 82.92 83.14 83.6
-LoRA 87.57 87.33 84.76 77.58 77.13 79.42 83.26 83.23 83.4 83.87
-CoACT (Ours) 93.52 93.12 90.66 82.07 82.75 84.1 86.77 86.64 86.91 86.86
-CI100 Prototype learning 95.8 88.95 84.83 82.75 80.1 77.6 77.1 76.46 76.68 75.58
-Linear tuning 95.87 89.02 84.88 82.78 80.11 77.6 77.11 76.47 76.69 75.58
-LoRA 96.37 89.63 85.36 83.39 80.51 77.94 77.37 76.71 76.92 75.81
-CoACT (Ours) 97.13 90.33 86.82 85.04 81.99 80.68 79.97 79.23 79.43 78.31
-CO211 Prototype learning 22.41 16.98 13.97 11.45 10.76 9.55 8.82 8.08 7.76 7.42
-Linear tuning 24.07 17.6 14.72 11.74 10.79 9.61 8.85 8.11 7.79 7.45
-LoRA 24.11 17.63 14.75 11.76 10.77 9.6 8.84 8.1 7.79 7.45
-CoACT (Ours) 23.64 16.43 14.05 11.43 10.56 9.43 8.75 8.02 7.72 7.42
-CUB200 Prototype learning 87.17 78.1 79.92 77.19 78.77 78.79 76.74 74.93 73.55 74.1
-Linear tuning 90.33 79.58 80.92 77.93 79.36 79.28 77.19 75.32 73.89 74.41
-LoRA 91.83 80.62 81.54 78.25 79.6 79.55 77.46 75.58 74.16 74.65
-CoACT (Ours) 94.78 82.38 83.26 80.35 81.56 81.13 79.58 78.13 77.11 77.38
-DTD Prototype learning 86.0 77.0 74.0 68.62 65.0 61.67 58.57 57.0 54.72
-Linear tuning 87.0 77.75 74.0 68.75 65.3 61.75 58.57 57.0 54.72
-LoRA 86.17 77.17 73.94 68.62 65.07 61.64 58.55 56.98 54.7
-CoACT (Ours) 88.17 80.25 75.11 69.08 65.27 62.08 58.9 57.4 55.11
-EUSAT Prototype learning 100.0 97.99 87.08 84.91 75.55 72.19 65.45 64.51 60.66 62.2
-Linear tuning 100.0 97.99 87.08 84.91 75.55 72.19 65.45 64.51 60.66 62.2
-LoRA 100.0 97.99 87.08 84.91 75.55 72.19 65.45 64.51 60.66 62.2
-CoACT (Ours) 100.0 98.1 87.4 85.05 75.7 72.51 65.56 64.5 60.74 62.25
-AirCr. Prototype learning 19.77 18.34 13.5 17.48 17.71 17.13 14.91 15.88 15.28 15.51
-Linear tuning 19.77 17.12 13.1 16.76 17.85 16.8 14.67 15.62 15.0 15.32
-LoRA 19.52 16.84 13.0 16.69 17.79 16.75 14.67 15.62 15.0 15.32
-CoACT (Ours) 28.57 15.45 12.57 19.85 20.7 21.08 19.42 19.77 19.09 18.98
-Food101 Prototype learning 82.57 79.79 80.85 80.09 77.48 76.03 76.18 74.32 72.35 71.09
-Linear tuning 83.55 80.25 81.18 80.4 77.73 76.21 76.33 74.43 72.44 71.18
-LoRA 83.26 80.23 81.24 80.48 77.79 76.29 76.38 74.49 72.5 71.24
-CoACT (Ours) 83.77 80.75 81.62 80.82 78.08 76.64 76.69 74.81 72.81 71.59
-GTSRB Prototype learning 33.89 21.29 27.6 24.74 22.66 18.5 16.6 16.05 15.98 15.53
-Linear tuning 40.83 26.59 31.25 26.87 24.64 20.45 18.4 17.75 17.52 16.92
-LoRA 40.64 26.89 31.3 27.03 24.89 20.84 18.79 18.14 17.88 17.23
-CoACT (Ours) 54.76 26.72 39.12 37.74 33.05 29.57 27.49 26.96 26.09 25.11
-MiniIN Prototype learning 99.4 98.3 97.43 97.32 97.04 97.03 96.41 96.31 95.46 95.31
-Linear tuning 99.4 98.3 97.43 97.32 97.04 97.03 96.41 96.32 95.47 95.32
-LoRA 99.4 98.25 97.4 97.3 97.02 97.02 96.4 96.31 95.46 95.32
-CoACT (Ours) 99.47 98.32 97.46 97.33 97.03 97.03 96.42 96.31 95.48 95.34
-FLO102 Prototype learning 100.0 100.0 97.25 99.0 98.5 98.57 98.5 98.11 98.3 98.09
-Linear tuning 100.0 100.0 97.25 99.0 98.5 98.57 98.5 98.11 98.3 98.09
-LoRA 100.0 100.0 97.25 99.0 98.5 98.57 98.5 98.11 98.3 98.09
-CoACT (Ours) 100.0 100.0 97.25 98.93 98.44 98.52 98.46 98.22 98.4 98.18
-Pets Prototype learning 93.76 93.38 90.48 92.51 92.81 92.64 88.78 89.3 88.1
-Linear tuning 93.76 93.38 90.48 92.51 92.81 92.64 88.78 89.3 88.1
-LoRA 93.76 93.38 90.48 92.51 92.81 92.64 88.78 89.3 88.1
-CoACT (Ours) 95.31 94.65 91.34 93.04 93.19 93.0 89.12 89.51 87.66
-RES45 Prototype learning 78.84 77.07 70.43 63.98 58.49 54.5 51.73 51.71 49.83
-Linear tuning 78.84 77.07 70.43 63.98 58.49 54.53 51.73 51.71 49.83
-LoRA 79.21 77.56 70.66 64.02 58.54 54.53 51.72 51.73 49.82
-CoACT (Ours) 97.12 89.64 80.53 75.65 70.53 66.33 64.1 64.34 62.62
-Cars Prototype learning 45.67 29.35 26.89 25.17 23.69 21.65 20.19 19.94 18.9
-Linear tuning 49.06 29.35 26.75 25.47 23.89 21.94 20.27 20.04 18.96
-LoRA 49.09 29.63 26.9 25.47 23.91 21.95 20.28 20.09 18.94
-CoACT (Ours) 61.82 38.24 35.84 33.09 31.4 28.35 25.4 25.76 24.4
-SUN397 Prototype learning 86.95 83.75 78.98 75.5 72.83 70.52 68.65 66.61 65.68
-Linear tuning 88.72 84.63 79.54 75.89 73.17 70.8 68.89 66.84 65.87
-LoRA 89.5 84.9 79.76 76.01 73.24 70.9 68.94 66.93 65.9
-CoACT (Ours) 89.82 84.75 79.91 76.08 73.36 70.99 69.01 67.08 66.11
-VOC Prototype learning 98.86 94.27 92.37 90.44 88.46 81.79 81.28 65.39 64.4 64.72
-Linear tuning 98.86 94.27 92.37 90.44 88.46 81.79 81.28 65.39 64.4 64.72
-LoRA 98.86 94.29 92.38 90.45 88.46 81.79 81.28 65.39 64.4 64.72
-CoACT (Ours) 99.62 94.7 92.45 90.48 88.35 81.75 81.31 65.06 64.1 64.45
-Table 9: Comparison to prior works on traditional FSCIL on CIFAR-100 with 60 base classes and 5-way 5-shot incremental setting.
-Method PT. Backbone Acc. in each session (%) ↑
-0 1 2 3 4 5 6 7 8
-Rebalancing [16] 61.31 47.80 39.31 31.91 25.68 21.35 18.67 17.24 14.17
-iCaRL [37] 61.31 46.32 42.94 37.63 30.49 24.00 20.89 18.80 17.21
-TOPIC [43] 61.31 50.09 45.17 41.16 37.48 35.52 32.19 29.46 24.42
-IDLVQ-C [3] 64.77 59.87 55.93 52.62 49.88 47.55 44.83 43.14 41.84
-FSLL [31] 66.48 61.75 58.16 54.16 51.10 48.53 46.54 44.20 42.28
-FSLL+SS [31] 68.85 63.14 59.24 55.23 52.24 49.65 47.74 45.23 43.92
-F2M [41] 67.28 63.80 60.38 57.06 54.08 51.39 48.82 46.58 44.65
-CEC [48] 73.07 68.88 65.26 61.19 58.09 55.57 53.22 51.34 49.14
-MetaFSCIL [7] 74.50 70.10 66.84 62.77 59.48 56.52 54.36 52.56 49.97
-CLOM [53] 74.2 69.83 66.17 62.39 59.26 56.48 54.36 52.16 50.25
-C-FSCIL [15] 77.47 72.40 67.47 63.25 59.84 56.95 54.42 52.47 50.47
-LIMIT [51] 73.81 72.09 67.87 63.89 60.70 57.77 55.67 53.52 51.23
-FACT [50] 74.60 72.09 67.56 63.52 61.38 58.36 56.28 54.24 52.10
-SAVC [42] 79.85 73.70 69.37 65.28 61.91 59.27 57.24 54.97 53.12
-SoftNet [21] 79.88 75.54 71.64 67.47 64.45 61.09 59.07 57.29 55.33
-BOT [39] 80.25 77.20 75.09 70.82 67.83 64.86 62.73 60.52 58.75
-SV-T [35] SwinT 86.77 82.82 80.36 77.20 76.06 74.00 72.92 71.68 69.75
-CPE-CLIP[9] CLIP-B/16 87.83 85.86 84.93 82.85 82.64 82.42 82.27 81.44 80.52
-Ours ViT-B/16 90.46 88.46 88.11 86.94 86.98 86.52 86.39 86.0 84.63
-Table 10: Comparison to prior works across the base and incremental sessions on CUB-200.
-Method Acc. in each session (%) ↑
-0 1 2 3 4 5 6 7 8 9 10
-iCaRL [37] 68.68 52.65 48.61 44.16 36.62 29.52 27.83 26.26 24.01 23.89 21.16
-EEIL [2] 68.68 53.63 47.91 44.20 36.30 27.46 25.93 24.70 23.95 24.13 22.11
-TOPIC [43] 68.68 62.49 54.81 49.99 45.25 41.40 38.35 35.36 32.22 28.31 26.28
-Rebalancing [16] 68.68 57.12 44.21 28.78 26.71 25.66 24.62 21.52 20.12 20.06 19.87
-SPPR [52] 68.68 61.85 57.43 52.68 50.19 46.88 44.65 43.07 40.17 39.63 37.33
-MetaFSCIL[7] 75.90 72.41 68.78 64.78 62.96 59.99 58.30 56.85 54.78 53.82 52.64
-F2M [41] 81.07 78.16 75.57 72.89 70.86 68.17 67.01 65.26 63.36 61.76 60.26
-CEC [48] 75.85 71.94 68.50 63.50 62.43 58.27 57.73 55.81 54.83 53.52 52.28
-FACT [50] 75.90 73.23 70.84 66.13 65.56 62.15 61.74 59.83 58.41 57.89 56.94
-LIMIT [51] 75.89 73.55 71.99 68.14 67.42 63.61 62.40 61.35 59.91 58.66 57.41
-SoftNet [21] 78.07 74.58 71.37 67.54 65.37 62.60 61.07 59.37 57.53 57.21 56.75
-SAVC [42] 81.85 77.92 74.95 70.21 69.96 67.02 66.16 65.30 63.84 63.15 62.50
-BOT [39] 82.31 78.03 75.45 70.99 71.06 67.85 67.44 66.05 64.95 64.31 63.75
-CPE-CLIP [9] 81.58 78.52 76.68 71.86 71.52 70.23 67.66 66.52 65.09 64.47 64.60
-SV-T [35] 84.19 82.63 81.21 78.97 79.38 77.64 77.55 75.71 75.91 75.77 76.17
-Ours 88.68 86.26 85.83 83.38 83.52 81.71 81.77 81.77 81.02 80.76 81.19
-Table 11: Comparison to prior works across the base and incremental sessions on miniImageNet.
-Method Acc. in each session (%) ↑
-0 1 2 3 4 5 6 7 8
-iCaRL [37] 71.77 61.85 58.12 54.60 51.49 48.47 45.90 44.19 42.71
-Rebalancing [16] 72.30 66.37 61.00 56.93 53.31 49.93 46.47 44.13 42.19
-TOPIC [43] 61.31 50.09 45.17 41.16 37.48 35.52 32.19 29.46 24.42
-EEIL [2] 61.31 46.58 44.00 37.29 33.14 27.12 24.10 21.57 19.58
-FSLL [31] 66.48 61.75 58.16 54.16 51.10 48.53 46.54 44.20 42.28
-FSLL+SS [31] 68.85 63.14 59.24 55.23 52.24 49.65 47.74 45.23 43.92
-F2M [41] 72.05 67.47 63.16 59.70 56.71 53.77 51.11 49.21 47.84
-CEC [48] 72.00 66.83 62.97 59.43 56.70 53.73 51.19 49.24 47.63
-MetaFSCIL [7] 72.04 67.94 63.77 60.29 57.58 55.16 52.90 50.79 49.19
-C-FSCIL [15] 76.40 71.14 66.46 63.29 60.42 57.46 54.78 53.11 51.41
-FACT [50] 72.56 69.63 66.38 62.77 60.60 57.33 54.34 52.16 50.49
-CLOM [53] 73.08 68.09 64.16 60.41 57.41 54.29 51.54 49.37 48.00
-LIMIT [51] 72.32 68.47 64.30 60.78 57.95 55.07 52.70 50.72 49.19
-SoftNet [21] 79.77 75.08 70.59 66.93 64.00 61.00 57.81 55.81 54.68
-SAVC [42] 81.12 76.14 72.43 68.92 66.48 62.95 59.92 58.39 57.11
-BOT [39] 84.30 79.59 75.49 71.4 68.45 65.12 62.20 60.52 59.57
-CPE-CLIP [9] 90.23 89.56 87.42 86.80 86.51 85.08 83.43 83.38 82.77
-SV-T [35] 88.75 87.92 86.07 84.84 84.30 83.24 82.22 82.28 82.38
-Ours 97.63 97.55 97.09 97.02 97.0 96.58 96.3 96.29 96.24
-Table 12: Ablation study on all datasets.
-Asy. Cont.
-Cal. Tun.
-Doub. Con.
-Caltech101
-CIFAR100
-Country211
-CUB200
-DTD
-EuroSat
-FGVCAircraft
-Food101
-GTSRB
-MiniImageNet
-Flowers102
-OxfordPets
-Resisc-45
-StanfordCars
-SUN397
-VOC 2007
-Average
-86.86 78.31 7.42 77.38 55.11 62.25 18.98 71.59 26.05 95.34 98.18 87.79 62.62 24.4 66.11 64.45 61.43
-85.89 78.45 7.45 77.27 54.59 62.91 18.20 71.83 26.29 95.04 98.01 88.25 61.77 24.14 66.01 64.89 61.31
-86.01 78.96 7.46 77.2 54.82 62.31 18.91 71.47 26.05 95.34 98.15 87.92 61.31 24.07 65.98 64.59 61.28
-84.62 78.37 7.49 76.44 54.83 62.22 18.23 71.39 18.02 95.34 98.09 88.1 57.21 23.18 65.94 64.72 60.26
-83.47 75.58 7.42 74.1 54.72 62.2 15.51 71.09 15.53 95.31 98.09 88.1 49.83 18.9 65.68 64.72 58.77
-Figure 9: Study on different parameters of asynchronous contrastive tuning, including the number of training epochs and number of adapter layers
-Figure 10: Study on different parameters of controlled fine-tuning, including learning rate, learning rate factor, encoder fine-tuning start layer, and encoder fine-tuning start epoch.
-Figure 11: Study on incremental fine-tuning parameters, $\lambda$, and $\gamma$.